### v1

# Strict Subgoal Execution: Reliable Long-Horizon Planning in Hierarchical Reinforcement Learning

Abstract
Long-horizon goal-conditioned tasks pose fundamental challenges for reinforcement learning (RL), particularly when goals are distant and rewards are sparse. While hierarchical and graph-based methods offer partial solutions, they often suffer from subgoal infeasibility and inefficient planning. We introduce Strict Subgoal Execution (SSE), a graph-based hierarchical RL framework that enforces single-step subgoal reachability by structurally constraining high-level decision-making. To enhance exploration, SSE employs a decoupled exploration policy that systematically traverses underexplored regions of the goal space. Furthermore, a failure-aware path refinement, which refines graph-based planning by dynamically adjusting edge costs according to observed low-level success rates, thereby improving subgoal reliability. Experimental results across diverse long-horizon benchmarks demonstrate that SSE consistently outperforms existing goal-conditioned RL and hierarchical RL approaches in both efficiency and success rate.
1 Introduction
Recent advances in reinforcement learning (RL) have achieved impressive success across various domains [1, 2] . In many real-world applications, agents must achieve specific objectives, motivating the development of goal-conditioned RL (GCRL), where agents learn to reach a designated goal state provided by the environment [3, 4, 5]. Unlike conventional RL methods that rely on carefully designed reward functions, GCRL allows agents to directly pursue target states based on goal specifications. However, in sparse-reward and long-horizon environments, goals are often distant, making exploration difficult and hindering effective learning due to a lack of intermediate guidance.
To address this, hierarchical RL (HRL) decomposes the decision process into a high-level policy that selects subgoals and a low-level policy that executes actions to reach them [6, 7]. Since subgoals are typically closer and more attainable than the final goal, this structure facilitates learning in long-horizon settings. Nevertheless, HRL can still fail when the selected subgoals are too difficult for the low-level policy to reach reliably. To mitigate this issue, graph-based HRL methods have been proposed [8, 9, 10, 11, 12, 13]. These approaches construct a graph over the goal space, where nodes represent states or regions and edges denote feasible transitions. Subgoal selection is guided by shortest paths on the graph, improving success rates in complex tasks. However, even graph-based HRL remains limited in performance when the final goal lies far from the current state, as the high-level policy may require too many steps, resulting in unstable training and poor scalability.
While recent approaches have explored graph-based GCRL without hierarchical structures to mitigate the limitations of high-level policies, such methods often lack the flexibility to handle environments involving multiple goals or diverse reward signals, due to the absence of high-level reasoning [14, 15]. To address these limitations, we propose a novel graph-based HRL framework, Strict Subgoal Execution (SSE), which retains the advantages of high-level planning while enhancing robustness and improving long-horizon goal-reaching by enforcing subgoal completion within a single high-level decision step. Our contributions are summarized as follows:
  1. 1.
    HRL Framework with Single-Step Subgoal Execution: We introduce a hierarchical framework that guarantees each subgoal is reached within a single high-level decision, removing step limits and reducing the number of high-level actions required for long-horizon tasks.
  2. 2.
    Decoupled Exploration for Goal Space Coverage: A dedicated exploration policy is decoupled from the reward-driven high-level policy to guide the agent through underexplored regions of the goal space, improving spatial coverage and training efficiency.
  3. 3.
    Failure-Aware Path Refinement: To enhance subgoal reliability, we adjust graph edge costs based on low-level failure statistics, encouraging path planning to avoid unstable transitions and improving the robustness of subgoal execution.
By ensuring that any high-level subgoal in the map is reachable in a single step, our SSE achieves robust performance even when the final goal lies far across the map. The high-level layer enables sequential decision-making in more complex environments with multiple goals or intermediate objectives. Through extensive evaluation on diverse long-horizon, goal-conditioned tasks, our method consistently demonstrates superior success rates and sample efficiency compared to prior GCRL and HRL approaches, validating the effectiveness of the proposed framework.
2 Preliminaries
2.1 Universal Markov Decision Process and Goal-Conditioned RL
We consider a universal Markov decision process (UMDP) defined as a tuple $(\mathcal{S},\mathcal{G},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{G}$ is the goal space, $\mathcal{A}$ is the action space, $P$ is the transition dynamics, $R$ is the reward function, and $\gamma\in(0,1]$ is the discount factor [3]. At each time step $t$, the agent observes a goal $g\in\mathcal{G}$ and state $s_{t}$, selects an action $a_{t}\sim\pi(\cdot|s_{t},g)$, and receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1},g)$ and next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The goal of GCRL is to learn a goal-conditioned policy $\pi$ that maximizes the expected return $\sum_{t=0}^{H}r_{t}$, where $H$ is the episode length. When $\mathcal{S}\neq\mathcal{G}$, we assume the existence of a mapping $\phi$ such that $\phi(s)\in\mathcal{G}$, allowing the agent to infer goal progress from the current state. The goal $g$ may be fixed or randomly sampled per episode; while random goals aid generalization through diversity, the fixed-goal setting is more challenging due to its singular target.
2.2 HRL Frameworks and Graph-Based Subgoal Planning in GCRL
In goal-conditioned settings, HRL addresses long-horizon challenges by decomposing the policy into a high-level policy $\pi^{h}$ and a low-level policy $\pi^{l}$ [6, 7]. Every $k$ steps, $\pi^{h}(\cdot|s_{t},g)$ selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$, which $\pi^{l}(\cdot|s_{t},\tilde{g}_{t})$ attempts to reach using an auxiliary reward [16, 17, 18]. However, when subgoals are too distant, the low-level policy may fail to reach them within the given horizon, and distance-based penalties can hinder learning, especially under sparse rewards. To mitigate this, graph-based approaches construct a goal-space graph $G=(V,E)$, where $V$ denotes a set of landmark nodes and $E$ consists of edges weighted by the expected effort required to transition between nodes. Landmark nodes are typically selected either by applying farthest point sampling (FPS) to choose states that are maximally distant from one another [12, 19, 20], or by placing virtual goals uniformly from a predefined grid over the goal space [15]. In general, the edge cost is defined as
$d(v_{1}\rightarrow v_{2}):=\log_{\gamma}(1+(1-\gamma)Q^{G}(v_{1},v_{2},\pi^{l})),$ (1)
where $Q^{G}$ estimates the feasibility of reaching $v_{2}$ from $v_{1}$ under $\pi^{l}$, defined as the value function $Q^{l}$ of the low-level policy or a predefined estimator. Full details on $Q^{G}$ are provided in Appendix B.1.
In graph-based HRL, given a subgoal $\tilde{g}_{t}$, the shortest path from the current state’s embedding $\phi(s_{t})$ to $\tilde{g}_{t}$ is computed via Dijkstra’s algorithm [21], producing a sequence of intermediate waypoints $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$, where each $\mathrm{wp}_{i}\in V$. The low-level policy $\pi^{l}(\cdot|s_{t},\mathrm{wp}_{i})$ guides the agent through these waypoints in order: once $\mathrm{wp}_{i}$ is reached, the next target subgoal for $\pi^{l}$ is updated to $\mathrm{wp}_{i+1}$, continuing until $\tilde{g}_{t}$ is reached. The high-level policy is trained on transitions $(s_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k-1}r_{j},s_{t+k})$ in buffer $\mathcal{B}^{h}$, while the low-level policy $\pi^{l}$ is trained using $(s_{t},\mathrm{wp}_{i},a_{t},r^{l}_{t},s_{t+1})$ in buffer $\mathcal{B}^{l}$, where the low-level reward $r^{l}_{t}=-1$ if the agent has not reached $\mathrm{wp}_{i}$, and $r^{l}_{t}=0$ otherwise [12, 19, 20]. In contrast, certain approaches dispense with high-level subgoals and directly train the low-level policy to reach the final goal $g$, under the assumption that intermediate subgoals are unnecessary [14, 15].
3 Related work
Advances in Goal-Conditioned RL
GCRL refers to RL settings where the agent is explicitly conditioned on a goal [22, 23, 24]. Modern GCRL methods typically employ Universal Value Function Approximators (UVFA) [3], which facilitate generalization across different goals by learning goal-conditioned policies. A key challenge in GCRL is low data efficiency, as transitions collected under one goal often become unusable for others [25]. To address this, many approaches perform goal relabeling in the replay buffer, transforming failed experiences into successful ones for alternative goals [26, 27, 28, 29, 30]. Beyond random goal sampling, recent work emphasizes structured exploration of the goal space to further improve sample efficiency and avoid local optima. This includes methods that prioritize rare or novel states [31, 32], generate goals using curriculum learning [29, 33], or discover diverse and semantically meaningful goals through unsupervised learning [34, 35, 36].
Subgoal Structuring and Planning in Hierarchical RL
Hierarchical frameworks for GCRL introduces multi-level policy structures to address long-horizon tasks by decomposing them into temporally abstracted subgoals. Foundational work has explored option-based [6] and manager-worker [7] frameworks, demonstrating the effectiveness of hierarchical decomposition in complex decision-making [37, 38, 9, 39]. To improve training stability in off-policy settings, data relabeling techniques have been proposed [40], while subgoal reachability has been enhanced through adjacency-based penalties and adversarial training [4, 16]. Recent advances incorporate graph structures to represent the state space and improve exploration in sparse-reward environments [8, 9, 10, 11, 12, 13]. Other methods leverage graph-based planning to better align high-level decisions with low-level policy execution, mitigating mismatches in temporal abstraction [19].
Structured Graph-Based Approaches for Goal-Conditioned RL
Graph-based approaches in GCRL offer a structured representation of the goal space, enabling planning over discrete landmarks and improved navigation in sparse-reward environments [10, 11, 12]. This direction has evolved through integration with latent modeling [13], hybrid control [41], policy-driven graph construction [42], and hierarchical planning [19]. To improve exploration, recent methods propose frontier-based expansion [20], curriculum strategies for reachability-aware goal selection [14], and virtual subgoal generation for broader coverage [15], all contributing to more efficient learning under sparse rewards.
4 Methods
4.1 Motivation: Rethinking Subgoal Execution in Graph-Based Hierarchical RL
Figure 1: Agent trajectory comparison in goal space $\mathcal{G}$: (a) conventional HRL with $k$-step limit, (b) proposed SSE with multiple subgoals, and (c) SSE selecting $g$ as the initial subgoal. In (a), reachable range is limited by $k$ steps and the agent continues without fully reaching the subgoal. In contrast, SSE allows arbitrary steps but enforces strict subgoal completion. States are shown as $\phi(s_{t})\in\mathcal{G}$.
Graph-based HRL shows strong potential for long-horizon tasks, yet conventional subgoal execution faces key limitations, particularly in sparse-reward and fixed-goal settings. In this section, we revisit standard designs and motivate a new framework that addresses these issues while preserving the hierarchical structure. Fig. 1(a) illustrates how conventional graph-based HRL operates in a fixed-goal setting. When the graph $G$ contains well-distributed landmark nodes, the high-level policy selects a subgoal $\tilde{g}_{t}$ from the current state $s_{t}$, and the low-level policy follows a waypoint path to reach it within a fixed $k$-step horizon. If the subgoal lies beyond this range, the agent may fail and proceed to the next subgoal $\tilde{g}_{t+k}$, causing variation in $\phi(s_{t+k})$ and destabilizing high-level learning. To mitigate this, some prior methods introduce distance-based penalties to favor closer subgoals, but this design still presents several limitations. In sparse-reward, long-horizon tasks, reachability penalties can dominate the learning signal, hindering high-level optimization. Failures such as collisions or instability often yield inconsistent representations $\phi(s_{t+k})$, degrading high-level updates. Moreover, due to the limited range of the low-level policy, reaching distant goals typically requires many intermediate subgoals, leading to inefficient learning when rewards are infrequent.
To address these challenges, we propose the Strict Subgoal Execution (SSE) framework, which enforces that each subgoal be reached within a single high-level step. This removes fixed step limits and allows subgoals to be placed arbitrarily far from the current state. Fig. 1(b) illustrates this mechanism: at each high-level decision step $t_{i}$ (with $t_{0}=0$), the agent is assigned a subgoal $\tilde{g}_{t}$, and the episode continues only if it successfully reaches $\tilde{g}_{t}$ from $s_{t}$; otherwise, it terminates immediately with no reward. This strict rule fundamentally departs from prior approaches by eliminating transitions resulting from failed subgoal attempts. In doing so, SSE provides more consistent learning signals and stabilizes high-level training. By removing artificial constraints such as step horizons and distance-based penalties, SSE emphasizes reliable subgoal completion as a key learning signal. This leads to more coherent planning in complex, long-horizon environments.
To ensure robust subgoal execution, SSE incorporates two additional components. First, a decoupled exploration policy independently explores under-visited regions and enhances the low-level policy’s ability to reach diverse subgoals, including those that are previously unexplored. Second, a failure-aware path refinement strategy updates edge costs based on subgoal failure ratio rather than geometric distance, promoting more reliable planning. Fig. 1(c) shows the resulting behavior after training with SSE: starting from the initial state, the agent selects the final goal $g$ directly as the first subgoal $\tilde{g}_{0}=g$ and reaches it without relying on intermediate waypoints. This example highlights the expressiveness and reliability of SSE. The proposed framework also generalizes well to complex tasks involving multiple goals or intermediate interactions, often completing them in only a few high-level steps.
4.2 Strict Subgoal Execution with Decoupled Exploration
In this section, we propose the Strict Subgoal Execution framework for graph-based HRL to improve subgoal reachability, as discussed in Section 4.1. While adopting the same graph construction and low-level learning strategies as prior graph-based HRL methods introduced in Section 2, our approach introduces a new high-level policy that ensures subgoal completion without relying on fixed-step limits. In prior works, subgoals are typically selected using Gaussian policies with small noise, often combined with penalties to encourage reachable subgoals. Although suitable under step-limited low-level control, these policies tend to overemphasize nearby goals and thus hinder effective exploration across the goal space. In contrast, our objective is to learn a high-level policy capable of reaching any point on the map in a single decision. To this end, we define an $\epsilon$-greedy high-level policy $\pi^{h}$ as:
$\pi^{h}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}\tilde{g}_{\max,t}:=\arg\max_{\tilde{g}\in\mathcal{G}}Q^{h}(\tilde{s}_{t},\tilde{g})&\text{with probability }1-\epsilon,\\ \text{randomly selected }\tilde{g}_{\mathrm{rand}}\in\mathcal{G}&\text{with probability }\epsilon,\end{cases}$ (2)
where $\tilde{g}_{\max,t}$ is the maximum subgoal, $\tilde{s}_{t}=(s_{t},t/H)$ augments $s_{t}$ with normalized time $t/H$ to reflect episode progress, and $Q^{h}$ is the high-level value function trained using RL. This formulation enables $\pi^{h}$ to explore the goal space extensively, independent of the agent’s current location.
Similar to prior work, we construct a waypoint path $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ from the current state $s_{t}$ to the subgoal $\tilde{g}_{t}$ using the graph $G=(V,E)$. The low-level policy $\pi^{l}$ sequentially follows these waypoints with unlimited step execution until the subgoal is reached. Let $k_{t}$ denote the number of steps required to reach $\tilde{g}_{t}$ from $s_{t}$; if the agent fails, $k_{t}$ may extend up to the full episode length $H$. Based on this, the $i$-th high-level decision time step is defined as $t_{i}=t_{i-1}+k_{t_{i-1}}$ for $i\geq 1$ with $t_{0}=0$. Unlike previous methods that continue the episode regardless of subgoal completion, SSE terminates the episode if the subgoal is not successfully reached. A subgoal is considered reached if $||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda$, where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\lambda>0$ is a distance threshold. If this condition is met, the next subgoal is generated; otherwise, the episode ends with no reward. The corresponding high-level transition for training $\pi^{h}$ is:
$\begin{cases}(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})&\text{if }||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda,\\ (\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})&\text{otherwise},\end{cases}$ (3)
where $\tilde{s}_{T}=(s_{T},1)$ is the terminal state. Under this structure, only successful subgoal completions yield positive rewards, thereby encouraging reliable subgoal execution. To leverage partial successes, we adopt an idea inspired by hindsight experience replay (HER) [26]. If the agent fails to reach $\tilde{g}_{t}$ but reaches the final visited waypoint $\mathrm{wp}_{\mathrm{final}}$ in the waypoint path, we treat it as a virtual subgoal and record the following transition: $(\tilde{s}_{t},g,\mathrm{wp}_{\mathrm{final}},\sum_{j=t}^{t+k^{\prime}_{t}-1}r_{j},\tilde{s}_{t+k^{\prime}_{t}})$, where $k^{\prime}_{t}$ is the number of steps taken to reach $\mathrm{wp}_{\mathrm{final}}$, and it satisfies $||\phi(s_{t+k^{\prime}_{t}})-\mathrm{wp}_{\mathrm{final}}||<\lambda$. This allows the transition to be treated as successful and contributes positively to training. All successful and failed transitions are stored in a high-level replay buffer $\mathcal{B}^{h}$ and used to train $\pi^{h}$ and $Q^{h}$ via standard off-policy RL.
Figure 2: Initial subgoals at $t=0$ selected by $\pi^{h}$ and $\pi^{\mathrm{exp}}$, with corresponding Ant agent trajectories at (a) early, (b) intermediate, and (c) final training stages in the U-maze task. The goal space (agent positions in the map) is partitioned into grid cells $C_{\mathcal{G}}^{m}$. $\pi^{h}$ selects between $\tilde{g}_{\max}$ and $\tilde{g}_{\mathrm{rand}}$ to encourage broad coverage, while $\pi^{\mathrm{exp}}$ samples from $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, and $g$ to visit underexplored regions and the goal. Over time, unreachable areas are excluded from subgoal candidates via SSE.
While SSE encourages precise goal-directed behavior, its strict reward structure can limit exploration. To address this, we introduce a decoupled exploration policy $\pi^{\mathrm{exp}}$ that actively explores under-visited regions of the goal space. At each step $t$, $\pi^{\mathrm{exp}}$ selects a subgoal according to:
$\pi^{\mathrm{exp}}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}g&\text{with probability }\frac{1}{3},\\ \tilde{g}_{\max,t}&\text{with probability }\frac{1}{3},\\ \text{randomly selected }\tilde{g}_{\mathrm{novel}}\in C_{\mathcal{G}}^{m_{\mathrm{novel}}}&\text{with probability }\frac{1}{3},\end{cases}$ (4)
where $\tilde{g}_{\mathrm{novel}}$, referred to as the novel subgoal, is sampled from the least visited region $C_{\mathcal{G}}^{m_{\mathrm{novel}}}$ among $M$ grid cells $C_{\mathcal{G}}^{m}$ of size $d_{\mathcal{G}}$ that partition the goal space, $\mathcal{G}=\bigcup_{m=1}^{M}C_{\mathcal{G}}^{m}$. Here, $m_{\mathrm{novel}}$ is determined by $m_{\mathrm{novel}}=\arg\min_{m}N(C_{\mathcal{G}}^{m})$, where $N(C_{\mathcal{G}}^{m})$ denotes the number of episodes in which grid $C_{\mathcal{G}}^{m}$ has been visited. The policy $\pi^{\mathrm{exp}}$ follows the same SSE mechanism as $\pi^{h}$ and stores its transitions in the high-level buffer $\mathcal{B}^{h}$. This strategy promotes broad and effective exploration by encouraging visits to final goals, high-value subgoals, and novel subgoals in a balanced manner. Early in training, we rely solely on $\pi^{\mathrm{exp}}$ and gradually mix it with $\pi^{h}$ using a target ratio of $\eta:(1-\eta)$, where $\eta$ is the exploration control factor. By improving coverage of promising and under-explored regions without additional learning, $\pi^{\mathrm{exp}}$ enhances subgoal accessibility and high-level generalization.
To illustrate the behavior of the proposed method, Fig. 2 shows grid-wise visitation in the goal space and low-level trajectories toward various initial subgoals at three stages of training: (a) early (10K steps), (b) intermediate (150K steps), and (c) final (500K steps). The environment is a U-maze where a MuJoCo [43] Ant agent navigates to the final goal $g$ located at the upper left of the map. In (a), most of the goal space remains unexplored, and initial subgoals selected by $\pi^{\mathrm{exp}}$ and $\pi^{h}$, including the final goal, random subgoals, and novel regions, result in wide-ranging trajectories that promote broad exploration. As training progresses in (b), the agent expands its coverage across the map, and in (c), it consistently reaches the goal $g$. Notably, SSE enables the agent to reach any reachable subgoal in a single high-level step, regardless of distance, supporting reliable execution and efficient long-range planning. This allows the agent to solve complex tasks using only a few high-level decisions.
4.3 Failure-Aware Path Refinement
In the SSE framework, reliable subgoal execution is crucial since each subgoal must be reached within a single high-level step. As described in Section 2, graph-based methods use Dijkstra’s algorithm to compute waypoint paths from $\phi(s_{t})$ to $\tilde{g}_{t}$ on a graph $G=(V,E)$, with edge distances defined by $d$ in Eq. (1). However, this distance does not capture failure cases such as collisions or getting stuck, leading agents to often fail even along the shortest path. We observe that such failure-prone regions significantly hinder subgoal success. To address this, we introduce a failure-aware path refinement strategy that increases edge costs in unreliable regions, steering the planner toward safer alternatives.
Using the grid cells $C_{\mathcal{G}}^{m}$, we define $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ as the number of episodes where the agent fails to reach a subgoal located outside the current cell, i.e., $\tilde{g}_{t}\in C_{\mathcal{G}}^{m^{\prime}}$ with $m^{\prime}\neq m$, but the episode terminates within $C_{\mathcal{G}}^{m}$. This indicates failure to exit the current region. To discourage such behavior, we increase traversal costs in these areas by defining the failure ratio as $\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})=N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})/N(C_{\mathcal{G}}^{m})$, and refine the edge distance from node $v_{1}$ to $v_{2}$ as:
$\tilde{d}(v_{1}\rightarrow v_{2})=d(v_{1}\rightarrow v_{2})\times\max\left(1,c_{\mathrm{dist}}\cdot\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})\right),\quad\forall v_{2}\in C_{\mathcal{G}}^{m},\quad\forall m,$ (5)
where $d$ is the original edge distance defined in Eq. (1), and $c_{\mathrm{dist}}>1$ is a scaling factor. A higher failure ratio increases the adjusted distance, encouraging Dijkstra’s algorithm to avoid unreliable regions. If no alternative exists, the original path is still taken to preserve subgoal reachability.
Fig. 3 illustrates the effect of the proposed path refinement in a bottleneck environment. In (a), without refinement, the agent repeatedly follows the shortest path through a narrow corridor near a wall, often resulting in failure. In (b), with refinement applied, increased edge costs in high-failure regions steer Dijkstra’s algorithm toward safer detours. When no alternatives exist (e.g., in the bottleneck), the agent still passes through, preserving reachability. This demonstrates that the refinement enhances subgoal success while maintaining overall reachability, supporting more stable execution in complex tasks. In summary, the proposed SSE framework is illustrated in Fig. 4, with a condensed version of the algorithm provided in Algorithm 1. The full algorithm and implementation details, including the graph construction method and training procedures, are provided in Appendix B.
Figure 3: Comparison of agent trajectories (blue lines) in a map with a bottleneck: (a) without path refinement and (b) with the proposed path refinement. Green lines represent the shortest waypoint paths computed via Dijkstra’s algorithm, while red areas denote grid cells with high failure ratios, i.e., $\text{ratio}_{\text{fail}}(C^{m}_{\mathcal{G}})>0.05$.
Figure 4: The proposed SSE framework.
Initialize policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, and graph $G$
for each iteration do
       for each episode do
             for each high-level selection step do
                   Sample a subgoal $\tilde{g}_{t}$ from $\pi^{h}$ or $\pi^{\mathrm{exp}}$ with ratio $\eta:(1{-}\eta)$
                  Plan the waypoint path from $\phi(s_{t})$ to $\tilde{g}_{t}$
                  Roll out low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ $\forall i$
                  if $\|\phi(s_{t+k_{t}})-\tilde{g}_{t}\|<\lambda$ (success) then
                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})$
                  else
                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})$ and terminate
                   end if
                  $N_{\mathrm{fail}}{(C_{\mathcal{G}}^{m})}+=1$ if $\phi(s_{t+k_{t}})\in C_{\mathcal{G}}^{m}$
             end for
            $N(C_{\mathcal{G}}^{m})+=1$ for all visited grids $C_{\mathcal{G}}^{m}$
       end for
      Update $Q^{h}$, $Q^{l}$, $\pi^{h}$, $\pi^{l}$ via off-policy RL
end for
Algorithm 1 Strict Subgoal Execution (SSE)
5 Experiments
In this section, we evaluate the proposed SSE framework on 9 challenging long-horizon tasks, including 5 AntMaze environments (U-maze, $\pi$-maze, AntMazeComplex, AntMazeBottleneck, and AntMazeDoubleBottleneck). Among these, U-maze has a simple layout, AntMazeBottleneck features narrow corridors, and the others present more complex structures. We also assess 2 KeyChest tasks (AntKeyChest, AntDoubleKeyChest), where the agent must collect 1 or 2 keys before reaching the final goal, even though the keys are not explicitly defined as goals. Additionally, we evaluate 2 Reacher tasks (ReacherWall, ReacherWallDoubleGoal), where a 3D robot must navigate obstacles to reach one or two goals. KeyChest and DoubleGoal tasks require intermediate objectives, making them ideal for testing high-level planning. Fig. 5 provides visualizations, with further details presented in Appendix C. While the main experiments focus on fixed-goal settings, additional comparisons for random-goal setups are included in Appendix D.1.
Figure 5: Considered long-horizon environments: 5 AntMaze, 2 KeyChest, and 2 Reacher tasks
5.1 Performance Comparison
Figure 6: Performance comparison on various long-horizon environments
We compare SSE with a range of baselines, including both hierarchical RL and recent graph-based methods. Specifically, we evaluate 2 HRL approaches: HIRO [40], which improves sample efficiency via hindsight goal relabeling, and HRAC [16], which penalizes subgoal selection based on reachability. We also include 3 graph-based HRL methods: HIGL [12], which applies intrinsic penalties over a graph, DHRL [19], which finds a path and gives waypoints to low-level as a goal via graph palnning, and NGTE [20], which expands graphs using novelty to better address fixed-goal settings. Additionally, we consider 2 graph-based methods without explicit high-level policies: PIG [42], which integrates graph structures into imitation learning to skip redundant subgoal actions, and BEAG [15], which uses breadth exploration with imaginary landmarks for goal-reaching. SSE is evaluated with the best hyperparameter settings ($c_{\mathrm{dist}}$, $d_{\mathcal{G}}$, $\eta$) from ablation studies, while all baselines use author-provided implementations. Detailed descriptions of each algorithm, along with the hyperparameter configurations for our proposed method, are provided in Appendix C.
Fig. 6 shows success rate comparisons across all tasks, where solid lines indicate the mean over 5 seeds and shaded areas denote standard deviations. From the result, SSE consistently outperforms both graph-based and conventional HRL methods across all benchmarks. Conventional HRL methods (DHRL, HIRO, PIG, HRAC, and HIGL) rely on random goal sampling for exploration but often struggle with fixed-goal tasks, while occasionally succeeding in random-goal setups provided in Appendix D.1, highlighting the increased challenge of fixed goals due to limited exploration opportunities. In relatively simple environments such as U-maze, $\pi$-maze, and AntMazeComplex, baseline methods like BEAG and NGTE demonstrate reasonable performance, but SSE typically converges more quickly. In bottleneck environments, SSE further excels by using failure-aware path refinement to avoid unstable regions as shown in Fig. 3. In more complex tasks like KeyChest and ReacherWallDoubleGoal, which require reaching intermediate objectives, baseline methods largely fail. This is because conventional HRL suffers from long high-level horizons, and goal-centric methods without high-level decision-making, such as BEAG, cannot reason about intermediate targets. In contrast, SSE mitigates these issues by strictly enforcing subgoal completion, reducing high-level decision steps. These results highlight the robustness, efficiency, and generalization of the proposed SSE framework. For practical comparison, we also evaluate the computational complexity against major baselines in Appendix D.2. The results show that our method achieves lower complexity per iteration, demonstrating its superiority in terms of computational efficiency.
5.2 Further Analysis and Ablation Studies
Fig. 8 presents a trajectory analysis of the proposed SSE framework in the AntDoubleKeyChest environment, illustrating how the agent progressively explores the map and collects both keys and the final goal. In the early stage (a) ($t\approx 300\text{K}$), the agent expands map coverage by sampling diverse subgoals, similar to simpler environments. As training progresses, the agent visits increasingly more regions, allowing it to reach the first key within a single high-level step, as shown in (b) ($t\approx 1\text{M}$). In a more advanced stage (c) ($t\approx 1.5\text{M}$), it collects both keys in just two high-level steps. Eventually, as shown in (d) ($t\approx 3\text{M}$), the agent completes the entire task, including both keys and the goal, in only three high-level steps. These results demonstrate that SSE allows the agent to reach any location in the map with a single high-level decision. As a result, it can solve complex multi-goal tasks using a minimal number of high-level steps, which highlights the effectiveness of SSE in long-horizon environments that require sequential decision-making for the high-level policy.
Figure 7: Trajectory analysis for SSE subgoals $\tilde{g}_{\max}$ in AntDoubleKeyChest at: (a) early stage, (b) reaches first key, (c) collects both keys, (d) reaches goal after collecting both keys (task success).
(a) Component evaluation
(b) Distance scaling $c_{\mathrm{dist}}$
(c) Grid size $d_{\mathcal{G}}$
(d) Exploration ratio $\eta$
Figure 8: Ablation study on AntDoubleKeyChest environment
We conduct an ablation study on AntDoubleKeyChest to evaluate the contribution of each component in SSE and the impact of key hyperparameters, as shown in Fig. 8. For the component analysis, we consider three variants: (1) SSE without path refinement (PR), which disables the failure-aware cost adjustment in graph planning; (2) SSE without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, $g$, or $\pi^{\mathrm{exp}}$, where we remove each subgoal type individually or disable the exploration policy altogether; and (3) SSE with a fixed low-level step (FLS), which replaces strict subgoal enforcement with a fixed-step high-level policy as in prior HRL methods. All variants lead to degraded performance, with the fixed-step setting failing to learn entirely, emphasizing the importance of strict subgoal execution and exploratory diversity. For hyperparameter sensitivity, setting $c_{\mathrm{dist}}=5$ achieves the optimal balance between penalizing failure-prone paths and avoiding excessive detours. Smaller values have little effect, while larger ones result in inefficient routing. Grid resolution remains stable, but overly fine grids (e.g., $d_{\mathcal{G}}=1$) slow learning due to increased planning costs. For the exploration ratio $\eta$, a value of 0.2 proves optimal, as lower values limit coverage too early and higher values delay the convergence of $\pi^{h}$. Despite some sensitivity, SSE consistently outperforms all baselines, demonstrating the SSE’s superiority. Additional trajectory analyses and ablation studies in other key environments are presented in Appendix E, further demonstrating the superiority of the proposed method.
6 Limitation
While SSE demonstrates strong performance across various long-horizon goal-conditioned tasks, it is currently limited to settings where the agent’s objective is solely to reach a specified goal. This assumption restricts its applicability to more complex tasks that require sequential or composite actions, such as robotic manipulation involving grasping and relocating objects, where multiple skill primitives may be necessary. Addressing such tasks would require extending the current framework to support a broader set of low-level skills and a high-level policy capable of selecting among them, presenting a promising direction for future work. Additionally, SSE relies on discretizing the goal space into grid cells to perform failure-aware path refinement. This approach is practical for environments with low-dimensional goal spaces such as 2D or 3D, commonly found in navigation and locomotion tasks. However, in high-dimensional goal spaces, the number of partitions increases rapidly, making exploration inefficient and coverage more difficult. To address this limitation, more advanced partitioning techniques such as clustering based on learned goal representations or adaptive sampling strategies could be considered to improve scalability in complex domains.
7 Conclusion
In this paper, We proposed Strict Subgoal Execution (SSE), a graph-based HRL framework designed to improve reliability and efficiency in long-horizon, goal-conditioned tasks. By enforcing single-step subgoal reachability, SSE enables robust high-level planning and significantly reduces decision horizons. The introduction of failure-aware path refinement and a decoupled exploration policy further enhances subgoal reliability and map coverage. Extensive experiments demonstrate that SSE consistently outperforms existing HRL and graph-based methods across a range of complex tasks. These results highlight the framework’s effectiveness in enabling stable and generalizable behavior, making it a promising approach for scalable hierarchical control in various long-horizon tasks.
Appendix A Broader Impact
The proposed SSE framework enhances the practicality and robustness of HRL in sparse-reward, long-horizon environments by eliminating fixed subgoal intervals and integrating failure-aware planning. This design reduces the dependency on reward shaping and broadens the applicability of HRL to real-world domains such as robotics, autonomous navigation, and long-term task planning. Furthermore, the graph-based structure of SSE improves interpretability, making it suitable for systems demanding transparency and reliability. However, its effectiveness still depends on careful design and validation; in safety-critical applications, poorly defined goals or sudden environmental shifts could lead to suboptimal behaviors. Nevertheless, SSE does not introduce specific societal or ethical concerns beyond standard RL considerations and poses no additional risks when applied responsibly.
Appendix B Implementation Details
This section presents additional implementation details of the proposed SSE framework. The graph construction method for the proposed SSE is detailed in Appendix B.1, while the training dynamics and implementation specifics of the framework are described in Appendix B.2. Each subsection highlights the design motivations and practical considerations for each module or mechanism.
B.1 Graph Construction of SSE
In this section, we describe how existing methods define the graph $G=(V,E)$ with the landmark node set $V$ and the edge set $E$ with edge distances $d$, and how the proposed SSE framework constructs this graph.
Landmark Node Set $V$ Construction
Existing HRL methods such as DHRL [19] and NGTE [20] construct the landmark node set $V$ by selecting visited states during exploration. They employ the Farthest Point Sampling (FPS) algorithm to identify landmark nodes that are far apart from each other. This approach ensures that frequently visited regions are well-represented, as it builds $V$ based on actual agent trajectories. However, it is limited to visited states, meaning unexplored or infrequently visited areas cannot be selected as landmarks, slowing exploration in those regions. In contrast, BEAG [15] accelerates exploration by partitioning the goal space into a grid structure, creating landmark nodes at each grid intersection. This method allows for faster exploration by using virtual goal positions as landmarks, independent of visitation frequency. The structured grid layout enables more systematic and efficient exploration. To leverage this advantage, the proposed SSE framework adopts a grid-based strategy for landmark node selection inspired by BEAG. Given a 2D goal space of size $x\times y$ and a grid size of $d_{\mathcal{G}}$, the landmark set $V$ is constructed as follows:
$V=\{(i\cdot d_{\mathcal{G}},j\cdot d_{\mathcal{G}})\in\mathcal{G}\mid i=0,\cdots,\frac{x}{d_{\mathcal{G}}}-1,~{}j=0,\cdots,\frac{y}{d_{\mathcal{G}}}-1\}.$ (6)
For a 3D goal space, the landmark node set is defined as $V=\{(i\cdot d_{\mathcal{G}},j\cdot d_{\mathcal{G}},k\cdot d_{\mathcal{G}})\in\mathcal{G}\mid i=0,\cdots,\frac{x}{d_{\mathcal{G}}}-1,~{}j=0,\cdots,\frac{y}{d_{\mathcal{G}}}-1,~{}k=0,\cdots,\frac{z}{d_{\mathcal{G}}}-1\}$. By constructing landmark nodes in this grid-based manner, SSE achieves faster and more structured exploration compared to visitation-based methods, ensuring efficient path planning and reliable subgoal execution.
Definition of Edge Distance
Given the landmark node set $V$, the edge set $E$ is defined as the collection of distances $d(v_{1}\rightarrow v_{2})$ between any two nodes $v_{1},v_{2}\in V$. For previous graph-based RL methods, the edge distance $d$ is computed as $d(v_{1}\rightarrow v_{2}):=\log_{\gamma^{l}}\left(1+(1-\gamma^{l})Q^{G}(v_{1},v_{2},\pi^{l})\right),$ where $\gamma^{l}$ is the low-level discount factor used for training the low-level policy $\pi^{l}$, and $Q^{G}$ is the value function estimating the traversal cost from $v_{1}$ to $v_{2}$, as described in Section 2. Existing HRL methods like DHRL [19] and NGTE [20] directly use the low-level value function $Q^{l}$, which is trained with a step-based reward of $-1$, to define the distance as the expected number of steps required for the low-level policy to navigate from $v_{1}$ to $v_{2}$. Although this method reflects actual navigation costs, it is sensitive to instability during $Q^{l}$ training, resulting in fluctuating edge distances. In contrast, BEAG [15] measures the distance using the Euclidean norm $d_{E}=||v_{1}-v_{2}||$, providing a stable but less accurate representation of traversal costs. To leverage the strengths of both approaches, SSE defines the edge distance $d(v_{1}\rightarrow v_{2})$ by combining $Q^{l}$ and $d_{E}$ as follows:
$d(v_{1}\rightarrow v_{2})=\frac{1}{2}\left[\log_{\gamma^{l}}\left(1+(1-\gamma^{l})Q^{l}(v_{1},v_{2},\pi^{l})\right)+\log_{\gamma^{l}}\left(1+(1-\gamma^{l})d_{E}(v_{1},v_{2})\right)\right].$ (7)
This hybrid formulation allows SSE to benefit from the stability of Euclidean distances when $Q^{l}$ is not fully converged, while still capturing the true traversal cost as $Q^{l}$ improves. As a result, the edge set $E$ is constructed as $E=\{d(v_{1}\rightarrow v_{2})\mid v_{1},v_{2}\in V\}$. Here, $d$ represents the raw edge distance before failure-aware path refinement is applied, ensuring both stability and adaptive accuracy in path estimation.
B.2 Detailed Loss Functions and Implementation of the SSE Framework
As described in Section 4, the proposed SSE framework is an HRL structure that employs a high-level policy $\pi^{h}$, an exploration policy $\pi^{\textrm{exp}}$, and a low-level policy $\pi^{l}$. The exploration policy does not require separate parameterization for high-level actions, while $\pi^{h}$ and $\pi^{l}$ are parameterized by $\theta^{h}$ and $\theta^{l}$, respectively, and represented as $\pi^{h}_{\theta^{h}}$ and $\pi^{l}_{\theta^{l}}$. To evaluate these policies, SSE defines the parameterized high-level value function $Q^{h}_{\psi^{h}}$ and the low-level value function $Q^{l}_{\psi^{l}}$, where $\psi^{h}$ and $\psi^{l}$ are the respective parameters. As mentioned in Section 2, the high-level policy and value function are trained to maximize external rewards, while the low-level policy and value function are optimized to reach designated waypoints incrementally. To facilitate this, the low-level policy receives the following reward $r^{l}_{t}$ at each time step:
$r^{l}_{t}=\begin{cases}0&\text{if }\|\phi(s_{t+1})-\mathrm{wp}_{i}\|<0.5,\textrm{ (agent reaches the current waypoint)}\\ -1&\text{otherwise},\end{cases}$ (8)
where $\mathrm{wp}_{i}$ is the target waypoint at the current timestep $t$, and $(\mathrm{wp}_{1},\cdots,\mathrm{wp}_{n})$ represent the shortest path from $\phi(s_{t})$ to the subgoal $\tilde{g}_{t}$. The high-level and low-level policies, along with their value functions, are trained using the transitions stored in the high-level buffer $\mathcal{B}^{h}$ and the low-level buffer $\mathcal{B}^{l}$ through the TD3 algorithm [44], a standard off-policy RL method. The value function losses for high-level and low-level policies are defined as follows:
$\small\mathcal{L}_{Q^{h}}(\psi^{h})$ $=\mathbb{E}_{B^{h}}\left[\left(Q^{h}(\tilde{s}_{t},\tilde{g}_{t})-\left(r^{h}_{t}+\gamma^{h}\min_{i=1,2}Q^{h}_{\bar{\psi}^{h}_{i}}(\tilde{s}_{t+k_{t}},\pi^{h}_{\theta^{h}}(\tilde{s}_{t+k_{t}},g))\right)\right)^{2}\right]$
$\mathcal{L}_{Q^{l}}(\psi^{l})$ $=\mathbb{E}_{B^{l}}\left[\left(Q^{l}(s_{t},a_{t})-\left(r^{l}_{t}+\gamma^{l}\min_{i=1,2}Q^{l}_{\bar{\psi}^{l}_{i}}(s_{t+1},\pi^{l}_{\theta^{l}}(s_{t+1},\mathrm{wp}_{i}))\right)\right)^{2}\right],$ (9)
where $r^{h}_{t}=\sum_{j=t}^{t+k_{t}-1}r_{j}$ for a successful trajectory and $r^{h}_{t}=0$ for a failed trajectory, $\bar{\psi}^{h}$ and $\bar{\psi}^{l}$ are target network parameters updated via exponential moving average (EMA), and $\gamma^{h}$ and $\gamma^{l}$ are the discount factors for training the high-level and low-level policies, respectively. The actor losses for optimizing the policies are defined as follows:
$\mathcal{L}_{\pi^{h}}(\theta^{h})=-\mathbb{E}_{B^{h}}\left[Q^{h}_{\psi^{h}}(\tilde{s}_{t},\pi_{\theta^{h}}^{h}(\tilde{s}_{t},g))\right],\quad\mathcal{L}_{\pi^{l}}(\theta^{l})=-\mathbb{E}_{B^{l}}\left[Q^{l}_{\psi^{l}}(s_{t},\pi_{\theta^{l}}^{l}(s_{t},\mathrm{wp}_{i}))\right].$ (10)
The parameters are optimized using the Adam optimizer [45] to minimize the respective loss functions. SSE distinguishes between the high-level discount factor $\gamma^{h}$ and the low-level discount factor $\gamma^{l}$, setting $\gamma^{l}=0.99$ as typical in RL, and $\gamma^{h}=0.4$ to limit return propagation across high-level steps, encouraging shorter path optimization. In the initial stages of training, rewards are sparse. To address this, the high-level buffer $\mathcal{B}^{h}$ is divided evenly, with half storing successful trajectories and the other half storing trajectories with zero reward. This design improves learning signals from successful experiences. The exploration ratio, which controls the balance between the exploration policy and the high-level policy, starts at $1{:}0$ and decays by 0.05 per iteration until it reaches $\eta{:}(1{-}\eta)$, as outlined in Section 4. This scheduling promotes exploration initially and shifts the focus to high-level learning as training progresses. If the low-level agent fails to reach the high-level subgoal $\tilde{g}_{t}$, the trajectory is marked as failed. In cases where the agent becomes stuck, such as flipping over or hitting obstacles, its position may remain unchanged for long periods. To improve sample efficiency, if no movement is detected for 500 steps, the trajectory is classified as failed and the episode is terminated. The complete SSE algorithm is provided in Algorithm 2.
Input: Graph $G=(V,E)$, goal $g$, mapping function $\phi$, threshold $\lambda$, exploration ratio $\eta$
Initialize: Policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, buffers $B^{h}$, $B^{l}$, grid cells $C_{\mathcal{G}}$
for each iteration do
       for each episode do
             Select the behavior policy: $\pi\leftarrow\pi^{\mathrm{exp}}$ with probability $\eta$, otherwise $\pi\leftarrow\pi^{h}$;
            for each high-level selection step do
                   Sample a subgoal $\tilde{g}_{t}\sim\pi(\tilde{s}_{t},g)$
                  Plan the waypoint path $\mathrm{wp}_{1:n}$ from $\phi(s_{t})$ to $\tilde{g}_{t}$ using Dijkstra’s algorithm over $G$ with $\tilde{d}$
                  for $i=1,\cdots,n$ do
                         Roll out the low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ to reach the waypoint $\mathrm{wp}_{i}$ within $k_{t}$ steps
                        Store the $k_{t}$ transitions $(s_{t},\mathrm{wp}_{i},a_{t},r_{t},s_{t+1})$ into the low-level buffer $\mathcal{B}^{l}$
                        Compute the reward sum: $r_{\mathrm{sum}}=\sum_{j=t}^{t+k_{t}-1}r_{j}$
                        if $\|\phi(s_{t+1})-\tilde{g}_{t}\|<\lambda$ then
                               Success: Store the high-level transition $(\tilde{s},g,\tilde{g}_{t},r_{\mathrm{sum}},\tilde{s}_{t+k_{t}})$ into $B^{h}$
                        else
                               Failure: Store $(\tilde{s},g,\tilde{g}_{t},0,\tilde{s}_{T})$ into $B^{h}$
                              $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})+=1$ for $m$ s.t. $\phi(s_{t+k_{t}})\in C_{\mathcal{G}}^{m}$
                              Terminate the episode
                         end if
                   end for
             end for
            $N(C_{\mathcal{G}}^{m})+=1$ for all visited cells $C_{\mathcal{G}}^{m}$
       end for
      Update $\psi^{h}$, $\psi^{l}$, $\theta^{h}$, $\theta^{l}$ to minimize RL losses $\mathcal{L}_{Q^{h}}$, $\mathcal{L}_{Q^{l}}$, $\mathcal{L}_{\pi^{h}}$, $\mathcal{L}_{\pi^{l}}$, respectively
end for
Algorithm 2 Strict Subgoal Execution (SSE)
Appendix C Experimental Setup
Our proposed framework is designed to be modular and general, enabling integration with a wide range of baseline methods. For comparison, we employ the official codebases provided by the original authors for HIRO, HRAC, HIGL, DHRL, NGTE, and BEAG. All baselines are run using the hyperparameters specified in their respective publications, and conducted on an NVIDIA RTX 3090 GPU with an Intel Xeon Gold 6348 CPU (Ubuntu 20.04). Appendix C.1 provides descriptions of the baseline algorithms along with links to their official code repositories. Appendix C.2 provides specifications of the nine environments shown in Fig.9, including their action and observation spaces, goal configurations, and episode horizons. The SSE-specific hyperparameter configurations for each environment are summarized in AppendixC.3.
C.1 Details of Other Baselines
  • HIRO [40] introduces a hierarchical architecture with relabeling of high-level transitions to account for changing low-level policies, thereby improving off-policy sample efficiency and stability. Open-source code of HIRO is available at https://github.com/watakandai/hiro_pytorch
  • HRAC [16] adds a learned adjacency constraint to ensure subgoal feasibility. It penalizes high-level selections that attempt transitions deemed unreachable within a limited horizon, thereby guiding the agent to learn feasible subgoal structures. Open-source code of HRAC is available at https://github.com/trzhang0116/HRAC
  • HIGL [12] incorporates a coverage-driven and novelty-driven landmark selection strategy. It performs graph-based planning via shortest paths and uses adjacency rewards to guide learning toward under-explored regions. Open-source code of HIGL is available at https://github.com/junsu-kim97/HIGL
  • DHRL [19] constructs a goal graph using Farthest Point Sampling and learns high-level behavior by planning over the graph. It emphasizes temporal abstraction and long-horizon planning via graph traversal and Q-learning. Open-source code of DHRL is available at https://github.com/jayLEE0301/dhrl_official
  • BEAG [15] employs value-function-driven imaginary landmarks to facilitate exploration of unvisited areas. It estimates landmark distances from a learned value function without relying solely on previously visited states, enabling efficient generalization. Open-source code of BEAG is available at https://github.com/ml-postech/BEAG
  • NGTE [20] drives novelty-based exploration by identifying frontier nodes (outposts) and prioritizing expansion toward less-visited regions of the goal graph, encouraging broad and diverse exploration. Open-source code of NGTE is available at https://github.com/ihatebroccoli/NGTE
C.2 Environmental Details
We follow standard benchmarks and configurations widely adopted in prior hierarchical reinforcement learning studies [19, 20, 15], and introduce several new environments designed to evaluate high-level decision-making capabilities such as AntKeyChest or AntDoubleKeyChest. The environments used in our experiments are visualized in Fig. 9, and their characteristics are described in Table 1.
Table 1: Summary of environment specifications.
Environment Spaces (Obs / Action) Goal space Reward Start / Goal Position Episode length
U-Maze Obs: 29-Dof
Action: 8-Dof
[-4,20]$\times$
[-4,20]
$1$ if goal reached
else 0
Start: (0,0)
Goal: (0,16)
600
$\pi$-Maze Obs: 29-Dof
Action: 8-Dof
[-4,36]$\times$
[-4,36]
$1$ if goal reached
else 0
Start: (8,0)
Goal: (24,0)
1000
AntMazeComplex Obs: 29-Dof
Action: 8-Dof
[-4,52]$\times$
[-4,52]
$1$ if goal reached
else 0
Start: (0,0)
Goal: (40,0)
2000
AntMazeBottle-
neck
Obs: 29-Dof
Action: 8-Dof
[-4,20]$\times$
[-4,20]
$1$ if goal reached
else 0
Start: (0,0)
Goal: (0,16)
600
AntMazeDouble-
Bottleneck
Obs: 29-Dof
Action: 8-Dof
[-4,20]$\times$
[-4,36]
$1$ if goal reached
else 0
Start: (0,0)
Goal: (16,32)
1200
AntKeyChest Obs: 30-Dof
Action: 8-Dof
[-4,36]$\times$
[-4,36]
$1$ if key reached
$5$ if goal reached
with key
Start: (0,0)
Key: (0,16)
Goal: (0,32)
2000
AntDouble-
KeyChest
Obs: 31-Dof
Action: 8-Dof
[-4,36]$\times$
[-4,36]
$1$ if key1 reached
$1$ if key2 reached
with key1
$5$ if goal reached
with two keys
Start: (0,0)
Key1: (16, 32)
Key2: (16, 0)
Goal: (32, 0)
3000
ReacherWall Obs: 17-Dof
Action: 7-Dof
[-1,1]$\times$
[-1,1]$\times$
[-1,1]
1 if goal reached Start:
(0.99, --0.19, 0)
Goal:
(0.6, 0.6, --0.1)
100
ReacherWall-
DoubleGoal
Obs: 22-Dof
Action: 7-Dof
[-1,1]$\times$
[-1,1]$\times$
[-1,1]
1 if goal reached
5 if both goals
reached
Start:
(0.99, --0.19, 0)
Goal1:
(0.4, 0.4, --0.1)
Goal2:
(0.4, --0.8, --0.1)
200
Figure 9: Considered long-horizon environments: 5 AntMaze, 2 KeyChest, and 2 Reacher tasks
AntKeyChest and AntDoubleKeyChest include key flags that are toggled from 0 to 1 when the agent successfully reaches the corresponding key location under a predefined condition. ReacherWallDoubleGoal contains two distinct goal positions, and two goal vectors are provided along with corresponding success flags, which are set to 1 upon reaching each goal. The success threshold for determining whether a target is reached is 5 in the AntMaze environments and 0.25 in the Reacher environments.
C.3 Hyperparameter Setup
Table 2 summarizes the common hyperparameters used across all environments. These configurations are based on a combination of parameter tuning and default settings from baseline implementations. In particular, the buffer size, batch size, and network architecture follow the original baseline code. Learning rates for each policy level, as well as the discount factors, were determined through a structured parameter search.
Table 3 presents environment-specific hyperparameters, including the minimum epsilon threshold $\epsilon_{\text{min}}$, the path refinement scaling factor $c_{\mathrm{dist}}$, the subgoal success threshold $\lambda$, and the grid resolution $d_{\mathcal{G}}$. These values were selected based on targeted parameter searches conducted for each environment.
In the case of AntDoubleKeyChest, a higher value of $\epsilon_{\text{min}}$ is required compared to other environments. This is because the task involves discovering multiple intermediate objectives, which increases the need for broad exploration. Other parameters such as the exploration ratio $\eta$ and the refinement scaling factor $c_{\mathrm{dist}}$ were also selected through environment-specific tuning, with further analyses reported in Appendix E.2.
Table 2: Common hyperparameter settings used in SSE.
Hyperparameter Value
Optimizer Adam
Replay buffer size 2,500,000
Batch size 1024
High-level actor learning rate $\alpha_{\mathrm{actor}}^{h}$ 0.000005
High-level critic learning rate $\alpha_{\mathrm{critic}}^{h}$ 0.00005
Low-level actor learning rate $\alpha_{\mathrm{actor}}^{l}$ 0.0001
Low-level critic learning rate $\alpha_{\mathrm{critic}}^{l}$ 0.001
High-level discount factor $\gamma^{h}$ 0.4
Low-level discount factor $\gamma^{l}$ 0.99
Target update rate $\tau$ 0.005
Table 3: Scenario-specific hyperparameters for SSE.
Scenario $\boldsymbol{\epsilon_{\text{min}}}$ $\boldsymbol{c_{\mathrm{dist}}}$ $\boldsymbol{\lambda}$ $\boldsymbol{d_{\mathcal{G}}}$ $\eta$
AntMaze
U-maze 0.1 5.0 2.0 2 0.1
AntMazeBottleneck 0.1 5.0 2.0 2 0.1
$\pi$-maze 0.1 5.0 2.0 2 0.2
AntMazeComplex 0.1 5.0 2.0 2 0.2
AntMazeDoubleBottleneck 0.1 10.0 2.0 2 0.1
AntKeyChest 0.1 5.0 2.0 2 0.2
AntDoubleKeyChest 0.2 5.0 2.0 2 0.2
Reacher
ReacherWall 0.1 5.0 0.15 1 0.2
ReacherWallDoubleGoal 0.1 5.0 0.15 1 0.2
Appendix D Additional Comparative Experiments
This section presents additional comparative experiments to further assess the generality and computational efficiency of the proposed SSE framework. We evaluate performance in a random-goal training setup in Appendix D.1, where goals are sampled uniformly from the entire reachable state space. Appendix D.2 evaluates computational characteristics, including convergence speed and per-episode compute time.
D.1 Performance Comparison under Random Goal Setups
Figure 10: Performance comparison in random goal setting
In Fig. 10, we compare experimental results under a random goal setting, where the goal is sampled uniformly from the entire valid state space during training, unlike the fixed-goal scenarios discussed in the main text. The reward signal remains sparse, posing a significant challenge for goal discovery and policy optimization. Graph-based methods such as BEAG, NGTE, and SSE, which can autonomously expand their subgoal graph during exploration, maintain strong performance even under random goal sampling, achieving success rates comparable to those in the fixed-goal setting. DHRL also benefits from the randomized goal distribution, particularly in simpler environments like U-maze and AntMazeBottleneck, where the training signal is more frequently encountered. On the other hand, methods such as HIRO, HIGL, and HRAC exhibit limited progress in most environments under random-goal conditions, due to their reliance on fixed-frequency subgoal selection, sparse reward setting, and lack of structured exploration mechanisms. In environments requiring multi-stage reasoning such as AntKeyChest, SSE is the only method that consistently discovers the key and solves the full task. NGTE, while hierarchical and exploration-driven, occasionally learns to acquire the key, but overall exhibits a very low success rate under this random goal configuration.
D.2 Computational Complexity Comparison
Table 4 reports the average per-episode computation time and the number of episodes required to reach a 60% success rate with NGTE and BEAG. We compare against NGTE and BEAG because they are the only baselines that succeed in at least one of the considered fixed-goal settings, making them the most relevant references for evaluating performance and sample efficiency in sparse reward environments. SSE achieves faster convergence across tasks due to its early termination of episodes upon subgoal failure and a reduced number of transitions per episode, which minimizes unnecessary computation and accelerates learning.
This advantage is particularly evident in long-horizon tasks such as AntKeyChest, where conventional HRL methods consume many timesteps even during failed attempts and require extended horizons for high-level planning. In contrast, in shorter-horizon scenarios such as AntMazeBottleneck, the computational benefits of SSE are less significant since subgoal transitions and failure terminations occur less frequently. The operations introduced by SSE, including the construction of $\tilde{s}$, evaluation of subgoal completion, and uniform sampling, are lightweight and have linear time complexity. As a result, SSE imposes minimal computational overhead while maintaining stable training dynamics.
Table 4: Average per-episode computation time (in seconds) and the number of episodes required to reach 60% success rate.
Scenario SSE BEAG NGTE
AntMazeBottleneck 5.94s 6.46s 12.65s
432 episodes 425.2 episodes 725.6 episodes
AntKeyChest 18.13s 57.29s 119.07s
1658 episodes fail fail
Appendix E Extended Analyses of the Proposed SSE Framework
To further validate the design and generality of SSE, this section presents extended analyses across several dimensions. Appendix E.1 provides qualitative trajectory visualizations that illustrate how SSE dynamically adapts its subgoal planning over time. Appendix E.2 presents ablation and sensitivity analyses to isolate the contribution of individual components and hyperparameters, including the path refinement scale $c_{\mathrm{dist}}$, grid resolution $d_{\mathcal{G}}$, and exploration ratio $\eta$.
E.1 Trajectory Analysis in Other Environments
To gain deeper insight into the behavior of SSE, we present trajectory analyses across representative environments.
Figure 11: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeBottleneck across: (a) early stage, (b) mid training, and (c) task success.
Fig. 11 illustrates learning progression in AntMazeBottleneck. At the early stage (a), the agent has not yet discovered feasible subgoals, and the high-level policy $\pi^{h}$ selects subgoals largely at random. During this phase, the exploration policy promotes coverage by gradually expanding into novel regions. In the mid stage (b), the agent fails to traverse the narrow bottleneck, primarily because the low-level policy $\pi^{l}$ lacks sufficient experience in that region and the associated failure statistics $\mathrm{ratio}_{\mathrm{fail}}$ remain underrepresented. By the final stage (c), SSE successfully leverages failure-aware refinement and subgoal selection, allowing $\pi^{h}$ to navigate through the bottleneck and reach the goal reliably.
Figure 12: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeDoubleBottleneck across: (a) early stage, (b) mid training, and (c) task success.
A similar trend is observed in AntMazeDoubleBottleneck as Fig. 12. Initially (a), most regions are unexplored, and subgoal selection remains uninformed. In the intermediate phase (b), the agent again struggles with the second bottleneck due to insufficient failure feedback. As training progresses (c), subgoals selected by $\pi^{h}$ become more effective, and the updated path refinement guides the agent through safer routes, enabling successful traversal of both bottlenecks and completion of the long-horizon task.
Figure 13: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntKeyChest across: (a) early stage, (b) after reaching the key, and (c) reaching the goal with the key (task success).
In Fig. 13, the agent in AntKeyChest begins by exploring the environment without a clear objective (a). Upon discovering the key (b), $\pi^{h}$ increasingly selects subgoals leading to the key location. In the final stage (c), the agent demonstrates the ability to sequentially reason over subtasks, first reaching the key and then navigating to the final goal with the required flag active, completing the task successfully.
E.2 Additional Ablation Studies
Component Evaluation
Figure 14: Component evaluation results in various maps
Fig. 14 presents the performance of SSE variants with key components ablated. Specifically, we evaluate the following variants: SSE without path refinement (PR), without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, or goal $g$ in $\pi^{\mathrm{exp}}$, without the entire $\pi^{\mathrm{exp}}$, and with a fixed low-level step size (FLS).
In complex tasks like AntKeyChest and AntDoubleKeyChest, removing max-value guidance ($\tilde{g}_{\max}$) or novelty-based exploration ($\tilde{g}_{\mathrm{novel}}$) leads to notable performance drops, due to poor goal-space coverage and delayed reward discovery. In contrast, simpler tasks like AntMazeBottleneck are less sensitive to these components. Path refinement similarly plays a larger role in long-horizon settings by mitigating unreliable transitions, while offering little benefit in short tasks with fewer failures. Removing the explicit goal signal (SSE w/o $g$) or exploration policy (SSE w/o $\pi^{\mathrm{exp}}$) consistently harms performance across all environments, confirming the importance of goal-conditioned reasoning and structured exploration. Lastly, replacing SSE’s flexible execution with a fixed-length scheduler (SSE w FLS) shows little effect in short tasks, but severely limits learning in long-horizon, sparse-reward environments like AntDoubleKeyChest.
Effect of Path Refinement Scaling Factor $c_{\mathrm{dist}}$
Figure 15: Distance scail $c_{\mathrm{dist}}$ analysis in various maps
Fig. 15 illustrates the effect of the path refinement scaling factor $c_{\mathrm{dist}}$ on performance across different environments. In short-horizon tasks such as AntMazeBottleneck, the sensitivity to $c_{\mathrm{dist}}$ is relatively low, and changes in its value have little impact on overall performance.
In contrast, for longer-horizon tasks such as AntKeyChest and AntDoubleKeyChest, the effect of $c_{\mathrm{dist}}$ becomes more significant. When the value increases from 1 to 5, the agent benefits from more reliable path selection due to moderate penalty scaling, which helps avoid failure-prone regions while maintaining reachability. However, setting $c_{\mathrm{dist}}$ to overly large values such as 10 or 20 leads to frequent detours and unstable planning, ultimately degrading both learning efficiency and task success.
Effect of Grid Size $d_{\mathcal{G}}$
Figure 16: Grid size $d_{\mathcal{G}}$ analysis in various maps
Fig. 16 illustrates how the grid resolution $d_{\mathcal{G}}$, used to partition the goal space for novelty identification, affects exploration. A coarse grid with $d_{\mathcal{G}}=4$ fails to capture fine-grained novelty, limiting the ability to distinguish under-visited regions. This shortcoming is pronounced in environments with narrow passages, such as AntMazeBottleneck, where structural bottlenecks are poorly reflected. In contrast, $d_{\mathcal{G}}=1$ creates overly fine partitions, fragmenting the goal space and introducing noise into failure statistics, which leads to unstable path updates and slower learning.
These results highlight the need for an appropriate balance between granularity and stability. Empirically, we use $d_{\mathcal{G}}=2$ for AntMaze environments to ensure reliable exploration while avoiding excessive updates. For Reacher environments, which feature a smaller and more structured goal space, a finer resolution of $d_{\mathcal{G}}=1$ is applied to retain sufficient detail without incurring instability.
Effect of Exploration Ratio $\eta$
Figure 17: Exploration ratio $\eta$ analysis in various maps
Fig. 17 presents the performance sensitivity with respect to the exploration ratio $\eta$. In short-horizon environments like AntMazeBottleneck, where high-level decisions are minimal and navigation is relatively simple, performance remains stable across a wide range of $\eta$ values. In such settings, large $\eta$ values are not necessary and may even hinder learning by reducing the frequency of value-guided subgoal selection, which slows the improvement of the high-level policy $\pi^{h}$.
In contrast, long-horizon environments such as AntDoubleKeyChest, which require intermediate decision-making and structured exploration, exhibit strong sensitivity to $\eta$. A small $\eta$ impedes the discovery of intermediate goals, while a large $\eta$ restricts exploitation of the learned value function $Q^{h}$ after critical behaviors are found. These results underscore the need to carefully tune $\eta$ to maintain a proper balance between exploration and exploitation, particularly in tasks with extended temporal dependencies.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the proposed framework (SSE), its key features and highlight consistent improvements across benchmark tasks in Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 6 describes limitations in high-dimensional goal spaces and in representing tasks that require abstract skills not decomposable into spatial subgoals.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work is empirical in nature and does not include formal theoretical claims that require proof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The proposed method is described in detail in Section 4, with full implementation details, including algorithmic procedures, training dynamics, and update rules, provided in Appendix B.2. These descriptions are sufficient to reproduce the main experimental results without requiring access to the source code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the full codebase and experiment scripts in the supplemental material, including detailed instructions for reproducing all main results. The code is organized for immediate execution with provided environment configurations.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All relevant training and evaluation details, including optimizer configurations, hyperparameter settings, and scenario-specific parameter choices, are thoroughly documented in Appendix C. These descriptions are sufficient to fully interpret and reproduce the experimental results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experimental results are averaged over 5 independent runs with different random seeds, and shaded regions in all figures represent the corresponding standard deviations, as detailed in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Details of the computational setup, including GPU types, memory, and per-run training time, are provided in Appendix C and Appendix D.2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research is based on standard benchmark environments, poses no foreseeable societal or ethical risks, and adheres fully to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Broader Impact is discussed in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve the release of models or data with potential for misuse or dual-use concerns.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All environments and baseline implementations are used under their official licenses and appropriately cited in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All introduced methods and environments are provided with detailed documentation to support reproducibility and understanding.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The research does not involve any human subjects or crowdsourcing components.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The research does not involve human subjects or any procedures requiring IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used in the development of the core methodology; usage was limited to writing and editing support.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 Long-horizon goal-conditioned tasks pose fundamental challenges for reinforcement learning (RL), particularly when goals are distant and rewards are sparse.
-While hierarchical and graph-based methods offer partial solutions, they often suffer from subgoal infeasibility and inefficient planning.
-We introduce Strict Subgoal Execution (SSE), a graph-based hierarchical RL framework that enforces single-step subgoal reachability by structurally constraining high-level decision-making.
-To enhance exploration, SSE employs a decoupled exploration policy that systematically traverses underexplored regions of the goal space.
-Furthermore, a failure-aware path refinement, which refines graph-based planning by dynamically adjusting edge costs according to observed low-level success rates, thereby improving subgoal reliability.
-Experimental results across diverse long-horizon benchmarks demonstrate that SSE consistently outperforms existing goal-conditioned RL and hierarchical RL approaches in both efficiency and success rate.
+While hierarchical and graph-based methods offer partial solutions, their reliance on conventional hindsight relabeling often fails to correct subgoal infeasibility, leading to inefficient high-level planning.
+To address this, we propose Strict Subgoal Execution (SSE), a graph-based hierarchical RL framework that integrates Frontier Experience Replay (FER) to separate unreachable from admissible subgoals and streamline high-level decision making.
+FER delineates the reachability frontier using failure and partial-success transitions, which identifies unreliable subgoals, increases subgoal reliability, and reduces unnecessary high-level decisions.
+Additionally, SSE employs a decoupled exploration policy to cover underexplored regions of the goal space and a path refinement that adjusts edge costs using observed low-level failures.
+Experimental results across diverse long-horizon benchmarks show that SSE consistently outperforms existing goal-conditioned and hierarchical RL methods in both efficiency and success rate.
+Our code is available at https://github.com/Jaebak1996/SSE
 1 Introduction
-Recent advances in reinforcement learning (RL) have achieved impressive success across various domains [1, 2] .
-In many real-world applications, agents must achieve specific objectives, motivating the development of goal-conditioned RL (GCRL), where agents learn to reach a designated goal state provided by the environment [3, 4, 5].
-Unlike conventional RL methods that rely on carefully designed reward functions, GCRL allows agents to directly pursue target states based on goal specifications.
+Recent advances in reinforcement learning (RL) have achieved impressive success across various domains (Mnih et al., 2013; Silver et al., 2016).
+Although prior works have enhanced sample efficiency through advanced exploration (Burda et al., 2018; Han and Sung, 2021b; a; Jo et al., 2024) and experience replay mechanisms (Schaul et al., 2015b; Han and Sung, 2017), they often struggle to adapt to changing objectives without extensive retraining.
+Consequently, Goal-Conditioned RL (GCRL) was introduced to explicitly condition the policy on a target goal, allowing agents to directly pursue diverse states specified by the environment (Schaul et al., 2015a; Levy et al., 2017; Nasiriany et al., 2019).
 However, in sparse-reward and long-horizon environments, goals are often distant, making exploration difficult and hindering effective learning due to a lack of intermediate guidance.
-To address this, hierarchical RL (HRL) decomposes the decision process into a high-level policy that selects subgoals and a low-level policy that executes actions to reach them [6, 7].
-Since subgoals are typically closer and more attainable than the final goal, this structure facilitates learning in long-horizon settings.
-Nevertheless, HRL can still fail when the selected subgoals are too difficult for the low-level policy to reach reliably.
-To mitigate this issue, graph-based HRL methods have been proposed [8, 9, 10, 11, 12, 13].
-These approaches construct a graph over the goal space, where nodes represent states or regions and edges denote feasible transitions.
-Subgoal selection is guided by shortest paths on the graph, improving success rates in complex tasks.
-However, even graph-based HRL remains limited in performance when the final goal lies far from the current state, as the high-level policy may require too many steps, resulting in unstable training and poor scalability.
-While recent approaches have explored graph-based GCRL without hierarchical structures to mitigate the limitations of high-level policies, such methods often lack the flexibility to handle environments involving multiple goals or diverse reward signals, due to the absence of high-level reasoning [14, 15].
-To address these limitations, we propose a novel graph-based HRL framework, Strict Subgoal Execution (SSE), which retains the advantages of high-level planning while enhancing robustness and improving long-horizon goal-reaching by enforcing subgoal completion within a single high-level decision step.
+Hierarchical reinforcement learning (HRL) addresses long-horizon goal-conditioned tasks by decomposing control into a high-level policy that proposes subgoals or skills and a low-level policy that executes actions to reach them (Bacon et al., 2017; Vezhnevets et al., 2017).
+While this decomposition can ease learning by introducing more attainable intermediate objectives, performance can degrade when proposed subgoals are not reliably reachable by the low-level policy.
+To mitigate this issue, graph-based HRL methods construct a graph over the goal space where nodes represent states or regions and edges encode feasible transitions (Zhang et al., 2018; Nachum et al., 2018a; Huang et al., 2019; Eysenbach et al., 2019; Kim et al., 2021; Zhang et al., 2021).
+This structure enables subgoal selection via shortest-path planning.
+Despite these advances, scalability remains challenging when the final goal is far from the current state because high-level planning may require many decisions and lead to unstable training.
+To reduce reliance on high-level policies, recent work has explored graph-based GCRL without explicit hierarchical structures (Lee et al., 2023; Yoon et al., 2024).
+However, removing high-level reasoning can reduce flexibility in environments with multiple goals or heterogeneous reward signals.
+To overcome these limitations, we propose Strict Subgoal Execution (SSE), a graph-based HRL framework that improves long-horizon performance by explicitly characterizing the reachability boundary of candidate subgoals.
+SSE integrates Frontier Experience Replay (FER) to distinguish high-level outcomes into failure transitions and partial-success transitions, localizing where execution breaks down and how far it progresses toward a subgoal.
+This frontier-based separation enables the agent to identify unreliable subgoals, train policies to execute admissible subgoals more accurately, and reduce unnecessary high-level decisions.
+SSE also expands goal-space coverage by actively exploring undervisited graph regions and improves planning reliability by refining paths to avoid transitions with frequent failures.
+As a result, SSE solves a broader range of complex long-horizon tasks.
 Our contributions are summarized as follows:
   1. 1.
-    HRL Framework with Single-Step Subgoal Execution: We introduce a hierarchical framework that guarantees each subgoal is reached within a single high-level decision, removing step limits and reducing the number of high-level actions required for long-horizon tasks.
+    Frontier Experience Replay (FER) for SSE: We introduce FER, which delineates the reachability frontier using two high-level samples: failure transitions and partial-success transitions.
+By precisely localizing where attempts fail and how far progress extends, FER identifies unreliable subgoals, increases subgoal reliability and reduces unnecessary high-level decisions.
   2. 2.
-    Decoupled Exploration for Goal Space Coverage: A dedicated exploration policy is decoupled from the reward-driven high-level policy to guide the agent through underexplored regions of the goal space, improving spatial coverage and training efficiency.
+    Decoupled Exploration for Goal Space Coverage: A dedicated exploration policy is decoupled from the return-driven high-level policy to traverse underexplored regions of the goal space, improving coverage and sample efficiency.
   3. 3.
-    Failure-Aware Path Refinement: To enhance subgoal reliability, we adjust graph edge costs based on low-level failure statistics, encouraging path planning to avoid unstable transitions and improving the robustness of subgoal execution.
-By ensuring that any high-level subgoal in the map is reachable in a single step, our SSE achieves robust performance even when the final goal lies far across the map.
-The high-level layer enables sequential decision-making in more complex environments with multiple goals or intermediate objectives.
-Through extensive evaluation on diverse long-horizon, goal-conditioned tasks, our method consistently demonstrates superior success rates and sample efficiency compared to prior GCRL and HRL approaches, validating the effectiveness of the proposed framework.
+    Failure-Aware Path Refinement: To improve subgoal reliability, we adjust graph edge costs based on low-level failure statistics, encouraging path planning to avoid unstable transitions and strengthening subgoal execution.
+Through extensive evaluation on diverse long-horizon goal-conditioned tasks, our method demonstrates higher success rates and better sample efficiency than prior GCRL and HRL approaches, validating the effectiveness of the proposed SSE framework.
 2 Preliminaries
-2.1 Universal Markov Decision Process and Goal-Conditioned RL
-We consider a universal Markov decision process (UMDP) defined as a tuple $(\mathcal{S},\mathcal{G},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{G}$ is the goal space, $\mathcal{A}$ is the action space, $P$ is the transition dynamics, $R$ is the reward function, and $\gamma\in(0,1]$ is the discount factor [3].
-At each time step $t$, the agent observes a goal $g\in\mathcal{G}$ and state $s_{t}$, selects an action $a_{t}\sim\pi(\cdot|s_{t},g)$, and receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1},g)$ and next state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
-The goal of GCRL is to learn a goal-conditioned policy $\pi$ that maximizes the expected return $\sum_{t=0}^{H}r_{t}$, where $H$ is the episode length.
-When $\mathcal{S}\neq\mathcal{G}$, we assume the existence of a mapping $\phi$ such that $\phi(s)\in\mathcal{G}$, allowing the agent to infer goal progress from the current state.
-The goal $g$ may be fixed or randomly sampled per episode; while random goals aid generalization through diversity, the fixed-goal setting is more challenging due to its singular target.
+2.1 Universal MDP, Goal-Conditioned RL, and Goal Relabeling Techniques
+We consider a universal Markov decision process (UMDP) $(\mathcal{S},\mathcal{G},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$, $\mathcal{G}$, and $\mathcal{A}$ denote the state, goal, and action spaces, $P$ the transition dynamics, $R$ the reward function, and $\gamma\in(0,1]$ the discount factor Schaul et al.
+(2015a).
+At time step $t$, the agent observes $(s_{t},g)$, samples $a_{t}\sim\pi(\cdot|s_{t},g)$, and receives $r_{t}=R(s_{t},a_{t},s_{t+1},g)$ with $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
+The goal of GCRL is to learn a goal-conditioned policy $\pi$ that maximizes the expected return $\sum_{t=0}^{H}r_{t}$ over horizon $H$.
+When $\mathcal{S}\neq\mathcal{G}$, GCRL typically assumes a mapping $\phi:\mathcal{S}\rightarrow\mathcal{G}$ projecting states into the goal space.
+The goal $g$ may be fixed or sampled per episode.
+In long-horizon goal-conditioned settings, learning is often inefficient due to sparse positive signals.
+Goal relabeling methods such as Hindsight Experience Replay (HER) (Andrychowicz et al., 2017) improve sample efficiency by treating later achieved states as substitute goals.
+HER replaces the goal $g$ with a future achieved goal $g^{\prime}=\phi(s_{t^{\prime}})$ for some $t^{\prime}\geq t$ and recomputes the reward:
+$(s_{t},a_{t},r_{t},s_{t+1},g)\ \mapsto\ (s_{t},a_{t},R(s_{t},a_{t},s_{t+1},g^{\prime}),s_{t+1},g^{\prime}).$
+This converts unsuccessful attempts into useful signals while the original transitions unchanged.
 2.2 HRL Frameworks and Graph-Based Subgoal Planning in GCRL
-In goal-conditioned settings, HRL addresses long-horizon challenges by decomposing the policy into a high-level policy $\pi^{h}$ and a low-level policy $\pi^{l}$ [6, 7].
-Every $k$ steps, $\pi^{h}(\cdot|s_{t},g)$ selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$, which $\pi^{l}(\cdot|s_{t},\tilde{g}_{t})$ attempts to reach using an auxiliary reward [16, 17, 18].
-However, when subgoals are too distant, the low-level policy may fail to reach them within the given horizon, and distance-based penalties can hinder learning, especially under sparse rewards.
-To mitigate this, graph-based approaches construct a goal-space graph $G=(V,E)$, where $V$ denotes a set of landmark nodes and $E$ consists of edges weighted by the expected effort required to transition between nodes.
-Landmark nodes are typically selected either by applying farthest point sampling (FPS) to choose states that are maximally distant from one another [12, 19, 20], or by placing virtual goals uniformly from a predefined grid over the goal space [15].
+In goal-conditioned settings, HRL addresses long-horizon challenges by decomposing control into a high-level policy $\pi^{h}$ and a low-level policy $\pi^{l}$ Bacon et al.
+(2017); Vezhnevets et al.
+(2017).
+In general, HRL frameworks in the GCRL domain typically assume that the goal space $\mathcal{G}$ is known to facilitate the high-level selection of feasible subgoals.
+Every $k$ steps, $\pi^{h}(\cdot\mid s_{t},g)$ selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$, which $\pi^{l}(\cdot\mid s_{t},\tilde{g}_{t})$ attempts to reach using an auxiliary reward  (Zhang et al., 2020; Pateria et al., 2021; Hutsebaut-Buysse et al., 2022).
+However, when subgoals are too distant, the low-level policy may fail to reach them within the given horizon, and distance-based penalties can hinder learning under sparse rewards.
+To mitigate this, graph-based approaches construct a goal-space graph $G=(V,E)$, where $V$ is a set of landmark nodes and $E$ contains edges weighted by the effort to transition between nodes.
+Landmarks are commonly chosen via farthest point sampling (FPS) (Kim et al., 2021; Lee et al., 2022) or placed on a predefined grid over the goal space (Yoon et al., 2024), assuming that the goal space is known to define the graph and the high-level policy.
 In general, the edge cost is defined as
@@
 where $Q^{G}$ estimates the feasibility of reaching $v_{2}$ from $v_{1}$ under $\pi^{l}$, defined as the value function $Q^{l}$ of the low-level policy or a predefined estimator.
-Full details on $Q^{G}$ are provided in Appendix B.1.
-In graph-based HRL, given a subgoal $\tilde{g}_{t}$, the shortest path from the current state’s embedding $\phi(s_{t})$ to $\tilde{g}_{t}$ is computed via Dijkstra’s algorithm [21], producing a sequence of intermediate waypoints $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$, where each $\mathrm{wp}_{i}\in V$.
-The low-level policy $\pi^{l}(\cdot|s_{t},\mathrm{wp}_{i})$ guides the agent through these waypoints in order: once $\mathrm{wp}_{i}$ is reached, the next target subgoal for $\pi^{l}$ is updated to $\mathrm{wp}_{i+1}$, continuing until $\tilde{g}_{t}$ is reached.
-The high-level policy is trained on transitions $(s_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k-1}r_{j},s_{t+k})$ in buffer $\mathcal{B}^{h}$, while the low-level policy $\pi^{l}$ is trained using $(s_{t},\mathrm{wp}_{i},a_{t},r^{l}_{t},s_{t+1})$ in buffer $\mathcal{B}^{l}$, where the low-level reward $r^{l}_{t}=-1$ if the agent has not reached $\mathrm{wp}_{i}$, and $r^{l}_{t}=0$ otherwise [12, 19, 20].
-In contrast, certain approaches dispense with high-level subgoals and directly train the low-level policy to reach the final goal $g$, under the assumption that intermediate subgoals are unnecessary [14, 15].
+In this paper, we set modified version of edge cost and details on $Q^{G}$ are provided in Appendix B.1.
+Given a subgoal $\tilde{g}_{t}$, graph-based HRL computes the shortest path from $\phi(s_{t})$ to $\tilde{g}_{t}$ via Dijkstra’s algorithm (Dijkstra, 1959), yielding waypoints $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ with $\mathrm{wp}_{i}\in V$.
+The low-level policy sequentially targets these waypoints, switching to $\mathrm{wp}_{i+1}$ once $\mathrm{wp}_{i}$ is reached, until $\tilde{g}_{t}$ is attained.
+The high-level policy is trained on transitions $(s_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k-1}r_{j},s_{t+k})$ in $\mathcal{B}_{F}^{h}$, while $\pi^{l}$ is trained on $(s_{t},\mathrm{wp}_{i},a_{t},r^{l}_{t},s_{t+1})$ in $\mathcal{B}^{l}$.
+Following prior work (Kim et al., 2021; Lee et al., 2022), the low-level reward is $r^{l}_{t}=0$ if $\phi(s_{t+1})-\mathrm{wp}_{i}||<\lambda$ and 1-1 otherwise, where $\lambda>0$ is a subgoal reachability threshold and $\phi$ is known.
+In contrast, some approaches omit high-level subgoals and directly train the low-level policy to reach the final goal $g$, assuming intermediate subgoals are unnecessary (Lee et al., 2023; Yoon et al., 2024).
 3 Related work
-Advances in Goal-Conditioned RL
-GCRL refers to RL settings where the agent is explicitly conditioned on a goal [22, 23, 24].
-Modern GCRL methods typically employ Universal Value Function Approximators (UVFA) [3], which facilitate generalization across different goals by learning goal-conditioned policies.
-A key challenge in GCRL is low data efficiency, as transitions collected under one goal often become unusable for others [25].
-To address this, many approaches perform goal relabeling in the replay buffer, transforming failed experiences into successful ones for alternative goals [26, 27, 28, 29, 30].
-Beyond random goal sampling, recent work emphasizes structured exploration of the goal space to further improve sample efficiency and avoid local optima.
-This includes methods that prioritize rare or novel states [31, 32], generate goals using curriculum learning [29, 33], or discover diverse and semantically meaningful goals through unsupervised learning [34, 35, 36].
-Subgoal Structuring and Planning in Hierarchical RL
-Hierarchical frameworks for GCRL introduces multi-level policy structures to address long-horizon tasks by decomposing them into temporally abstracted subgoals.
-Foundational work has explored option-based [6] and manager-worker [7] frameworks, demonstrating the effectiveness of hierarchical decomposition in complex decision-making [37, 38, 9, 39].
-To improve training stability in off-policy settings, data relabeling techniques have been proposed [40], while subgoal reachability has been enhanced through adjacency-based penalties and adversarial training [4, 16].
-Recent advances incorporate graph structures to represent the state space and improve exploration in sparse-reward environments [8, 9, 10, 11, 12, 13].
-Other methods leverage graph-based planning to better align high-level decisions with low-level policy execution, mitigating mismatches in temporal abstraction [19].
-Structured Graph-Based Approaches for Goal-Conditioned RL
-Graph-based approaches in GCRL offer a structured representation of the goal space, enabling planning over discrete landmarks and improved navigation in sparse-reward environments [10, 11, 12].
-This direction has evolved through integration with latent modeling [13], hybrid control [41], policy-driven graph construction [42], and hierarchical planning [19].
-To improve exploration, recent methods propose frontier-based expansion [20], curriculum strategies for reachability-aware goal selection [14], and virtual subgoal generation for broader coverage [15], all contributing to more efficient learning under sparse rewards.
+Goal-Conditioned RL and Hierarchical Approaches
+GCRL refers to RL settings where the agent is explicitly conditioned on a goal (Kaelbling, 1993; Liu et al., 2022; Colas et al., 2022).
+Modern GCRL typically employs Universal Value Function Approximators (UVFA) (Schaul et al., 2015a) to generalize across goals.
+A central challenge is solving long-horizon tasks with sparse rewards, where exploration is difficult.
+To address this, HRL introduces multi-level policies that decompose complex tasks into temporally abstract subgoals (Vezhnevets et al., 2017; Nachum et al., 2018b) or reusable skills (Eysenbach et al., 2018; Lee et al., 2025), thereby reducing the effective planning horizon.
+The effectiveness of this decomposition has been demonstrated in diverse settings (Barto and Mahadevan, 2003; Kulkarni et al., 2016; Nachum et al., 2018a; 2019).
+Beyond hierarchical structures, other methods improve sample efficiency through structured exploration, such as prioritizing novel states (Warde-Farley et al., 2018; Pong et al., 2019) or discovering useful goals via unsupervised learning (Mendonca et al., 2021; Ecoffet et al., 2021; Chane-Sane et al., 2021).
+Relabeling and Guidance Techniques in GCRL
+In sparse-reward GCRL, data relabeling is central for improving sample efficiency.
+HER converts failed trajectories into successes by replacing the intended goal with a future achieved goal (Andrychowicz et al., 2017).
+Since uniform hindsight-goal sampling can be suboptimal, subsequent work explores curriculum-based relabeling (Fang et al., 2019), novelty- or priority-driven goal sampling (Zeng et al., 2023), and hierarchical relabeling that better matches low-level behavior (Nachum et al., 2018b).
+Orthogonal guidance approaches discourage infeasible subgoals, for example via learned adjacency constraints (Zhang et al., 2020) or adversarially generated goals that are challenging yet achievable (Levy et al., 2017).
+Graph-based Approaches in GCRL
+Graph-based approaches have been introduced to GCRL to provide a structured representation of the goal space, enabling planning over discrete landmarks and improving navigation in sparse-reward environments (Huang et al., 2019; Eysenbach et al., 2019; Kim et al., 2021).
+Early work used graph structures to represent the state space and to guide exploration (Zhang et al., 2018; Nachum et al., 2018a), and this direction has since evolved through integration with latent modeling (Zhang et al., 2021) and policy-driven graph construction (Kim et al., 2023).
+More recent advances focus either on aligning high-level decisions with low-level execution via graph-based planning (Lee et al., 2022) or on enhancing exploration with strategies such as frontier-based expansion (Hoang et al., 2021; Park et al., 2024), curriculum-based goal selection (Lee et al., 2023), and virtual subgoal generation for broader coverage (Yoon et al., 2024).
+In related works, frontier-based exploration refers to exploring near boundary states.
+In our work, we also use the term frontier, but it denotes the boundary that identifies failure points and separates achievable from unachievable subgoals, which is a distinct objective from frontier-based exploration.
 4 Methods
-4.1 Motivation: Rethinking Subgoal Execution in Graph-Based Hierarchical RL
-Figure 1: Agent trajectory comparison in goal space $\mathcal{G}$: (a) conventional HRL with $k$-step limit, (b) proposed SSE with multiple subgoals, and (c) SSE selecting $g$ as the initial subgoal.
-In (a), reachable range is limited by $k$ steps and the agent continues without fully reaching the subgoal.
-In contrast, SSE allows arbitrary steps but enforces strict subgoal completion.
-States are shown as $\phi(s_{t})\in\mathcal{G}$.
-Graph-based HRL shows strong potential for long-horizon tasks, yet conventional subgoal execution faces key limitations, particularly in sparse-reward and fixed-goal settings.
-In this section, we revisit standard designs and motivate a new framework that addresses these issues while preserving the hierarchical structure.
+4.1 Motivation: Rethinking Subgoal Execution in Graph-Based HRL
+Figure 1: Agent trajectories in goal space $\mathcal{G}$.
+(a) Conventional HRL with HER relabels intermediate states as subgoals without enforcing exact subgoal completion, which lengthens high-level trajectories.
+(b) SSE with FER enforces exact subgoal completion, increasing subgoal reliability and reducing unnecessary high-level decisions, thereby improving learning efficiency.
+(c) After training, SSE reaches $g$ with few high-level steps, here in a single step in single-goal settings even from distant starts.
+Agent locations are $\phi(s_{t})\in\mathcal{G}$ and $t_{i}$ is the $i$-th high-level step.
+In GCRL, many methods improve goal-reaching performance by adopting hindsight relabeling techniques such as HER (Andrychowicz et al., 2017), which treat intermediate states as virtual goals to provide additional training signals.
+Recent graph-based HRL methods typically apply this idea to both the low level and the high level.
+While this benefits the low level, it introduces a critical issue at the high level.
 Fig.
-1(a) illustrates how conventional graph-based HRL operates in a fixed-goal setting.
-When the graph $G$ contains well-distributed landmark nodes, the high-level policy selects a subgoal $\tilde{g}_{t}$ from the current state $s_{t}$, and the low-level policy follows a waypoint path to reach it within a fixed $k$-step horizon.
-If the subgoal lies beyond this range, the agent may fail and proceed to the next subgoal $\tilde{g}_{t+k}$, causing variation in $\phi(s_{t+k})$ and destabilizing high-level learning.
-To mitigate this, some prior methods introduce distance-based penalties to favor closer subgoals, but this design still presents several limitations.
-In sparse-reward, long-horizon tasks, reachability penalties can dominate the learning signal, hindering high-level optimization.
-Failures such as collisions or instability often yield inconsistent representations $\phi(s_{t+k})$, degrading high-level updates.
-Moreover, due to the limited range of the low-level policy, reaching distant goals typically requires many intermediate subgoals, leading to inefficient learning when rewards are infrequent.
-To address these challenges, we propose the Strict Subgoal Execution (SSE) framework, which enforces that each subgoal be reached within a single high-level step.
-This removes fixed step limits and allows subgoals to be placed arbitrarily far from the current state.
-Fig.
-1(b) illustrates this mechanism: at each high-level decision step $t_{i}$ (with $t_{0}=0$), the agent is assigned a subgoal $\tilde{g}_{t}$, and the episode continues only if it successfully reaches $\tilde{g}_{t}$ from $s_{t}$; otherwise, it terminates immediately with no reward.
-This strict rule fundamentally departs from prior approaches by eliminating transitions resulting from failed subgoal attempts.
-In doing so, SSE provides more consistent learning signals and stabilizes high-level training.
-By removing artificial constraints such as step horizons and distance-based penalties, SSE emphasizes reliable subgoal completion as a key learning signal.
-This leads to more coherent planning in complex, long-horizon environments.
-To ensure robust subgoal execution, SSE incorporates two additional components.
-First, a decoupled exploration policy independently explores under-visited regions and enhances the low-level policy’s ability to reach diverse subgoals, including those that are previously unexplored.
-Second, a failure-aware path refinement strategy updates edge costs based on subgoal failure ratio rather than geometric distance, promoting more reliable planning.
-Fig.
-1(c) shows the resulting behavior after training with SSE: starting from the initial state, the agent selects the final goal $g$ directly as the first subgoal $\tilde{g}_{0}=g$ and reaches it without relying on intermediate waypoints.
-This example highlights the expressiveness and reliability of SSE.
-The proposed framework also generalizes well to complex tasks involving multiple goals or intermediate interactions, often completing them in only a few high-level steps.
-4.2 Strict Subgoal Execution with Decoupled Exploration
-In this section, we propose the Strict Subgoal Execution framework for graph-based HRL to improve subgoal reachability, as discussed in Section 4.1.
-While adopting the same graph construction and low-level learning strategies as prior graph-based HRL methods introduced in Section 2, our approach introduces a new high-level policy that ensures subgoal completion without relying on fixed-step limits.
-In prior works, subgoals are typically selected using Gaussian policies with small noise, often combined with penalties to encourage reachable subgoals.
-Although suitable under step-limited low-level control, these policies tend to overemphasize nearby goals and thus hinder effective exploration across the goal space.
-In contrast, our objective is to learn a high-level policy capable of reaching any point on the map in a single decision.
-To this end, we define an $\epsilon$-greedy high-level policy $\pi^{h}$ as:
-$\pi^{h}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}\tilde{g}_{\max,t}:=\arg\max_{\tilde{g}\in\mathcal{G}}Q^{h}(\tilde{s}_{t},\tilde{g})&\text{with probability }1-\epsilon,\\ \text{randomly selected }\tilde{g}_{\mathrm{rand}}\in\mathcal{G}&\text{with probability }\epsilon,\end{cases}$ (2)
-where $\tilde{g}_{\max,t}$ is the maximum subgoal, $\tilde{s}_{t}=(s_{t},t/H)$ augments $s_{t}$ with normalized time $t/H$ to reflect episode progress, and $Q^{h}$ is the high-level value function trained using RL.
-This formulation enables $\pi^{h}$ to explore the goal space extensively, independent of the agent’s current location.
-Similar to prior work, we construct a waypoint path $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ from the current state $s_{t}$ to the subgoal $\tilde{g}_{t}$ using the graph $G=(V,E)$.
-The low-level policy $\pi^{l}$ sequentially follows these waypoints with unlimited step execution until the subgoal is reached.
-Let $k_{t}$ denote the number of steps required to reach $\tilde{g}_{t}$ from $s_{t}$; if the agent fails, $k_{t}$ may extend up to the full episode length $H$.
-Based on this, the $i$-th high-level decision time step is defined as $t_{i}=t_{i-1}+k_{t_{i-1}}$ for $i\geq 1$ with $t_{0}=0$.
-Unlike previous methods that continue the episode regardless of subgoal completion, SSE terminates the episode if the subgoal is not successfully reached.
-A subgoal is considered reached if $||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda$, where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\lambda>0$ is a distance threshold.
-If this condition is met, the next subgoal is generated; otherwise, the episode ends with no reward.
-The corresponding high-level transition for training $\pi^{h}$ is:
-$\begin{cases}(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})&\text{if }||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda,\\ (\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})&\text{otherwise},\end{cases}$ (3)
-where $\tilde{s}_{T}=(s_{T},1)$ is the terminal state.
-Under this structure, only successful subgoal completions yield positive rewards, thereby encouraging reliable subgoal execution.
-To leverage partial successes, we adopt an idea inspired by hindsight experience replay (HER) [26].
-If the agent fails to reach $\tilde{g}_{t}$ but reaches the final visited waypoint $\mathrm{wp}_{\mathrm{final}}$ in the waypoint path, we treat it as a virtual subgoal and record the following transition: $(\tilde{s}_{t},g,\mathrm{wp}_{\mathrm{final}},\sum_{j=t}^{t+k^{\prime}_{t}-1}r_{j},\tilde{s}_{t+k^{\prime}_{t}})$, where $k^{\prime}_{t}$ is the number of steps taken to reach $\mathrm{wp}_{\mathrm{final}}$, and it satisfies $||\phi(s_{t+k^{\prime}_{t}})-\mathrm{wp}_{\mathrm{final}}||<\lambda$.
-This allows the transition to be treated as successful and contributes positively to training.
-All successful and failed transitions are stored in a high-level replay buffer $\mathcal{B}^{h}$ and used to train $\pi^{h}$ and $Q^{h}$ via standard off-policy RL.
+1(a) illustrates how graph-based HRL with HER operates: the high-level policy selects a subgoal $\tilde{g}_{t}$ given the current state $s_{t}$, and the low-level policy follows a waypoint path to reach it.
+If $\tilde{g}_{t}$ is unreachable due to limited skill, obstacles, or excessive distance, the agent stops at an intermediate state, and such subgoals should be avoided.
+When HER is applied, every visited state along the failed trajectory is treated as if it were the intended subgoal.
+This causes the high-level policy to repeatedly select ineffective subgoals and to waste many decision steps.
+In addition, the resulting transitions can vary widely for the same subgoal, which hinders stable learning.
+As shown in Fig.
+1(a), this produces unnecessarily long high-level trajectories, and even if the agent reaches the goal $g$, credit is spread over too many steps, preventing earlier decisions from being reinforced and often leading to failure on long-horizon tasks.
+To address this issue, we propose the Strict Subgoal Execution (SSE) framework, which updates the high-level policy with positive returns only when the low level successfully reaches the assigned subgoal.
+We instantiate this principle with Frontier Experience Replay (FER).
+Unlike relabeling methods such as HER that synthesize additional successes, FER delineates the reachability frontier by recording two types of high-level samples, failure transitions and partial-success transitions.
+By precisely localizing where attempts fail and how far progress extends, FER identifies unreliable subgoals, provides consistent training signals, and reduces unnecessary high-level decisions.
+As shown in Fig.
+1(b), this separation of success and failure keeps the resulting state $\phi(s_{t^{\prime}})$ closely aligned with the selected subgoal $\tilde{g}_{t}$, producing consistent high-level transitions and eliminating wasteful actions.
+Consequently, as illustrated in Fig.
+1(c), the resulting high-level policy solves tasks with far fewer decisions, often reaching the final goal $g$ in a single step in simple settings and handling multi-goal or long-horizon environments with only a few well-chosen subgoal selections.
+4.2 Strict Subgoal Execution with Frontier Experience Replay
+In this section, we describe the details of the SSE framework.
+We first define FER, the key component of SSE, which marks the reachability frontier by recording two high-level sample types in addition to standard successes: failure transitions that stop at the point of failure and partial-success transitions that record the last reliably reached waypoint.
+To formalize this, we basically follow the HRL setup from Section 2: At each time $t$, the policy selects a subgoal $\tilde{g}_{t}\in\mathcal{G}$ under the common assumption that the goal space $\mathcal{G}$ is known.
+A waypoint path $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ on the graph $G=(V,E)$ is then generated, and the low-level policy $\pi^{l}$ follows this path until termination at time $t^{\prime}$.
+We regard the subgoal as reachable if $\|\phi(s_{t^{\prime}})-\tilde{g}_{t}\|<\lambda$; otherwise the attempt is treated as a failure.
+In the failure case, $\mathrm{wp}_{\mathrm{final}}$ denotes the last waypoint reached within tolerance before failure.
+Based on this notion of reachability, we define FER as follows.
+Definition 4.1 (Frontier Experience Replay)
+The high-level replay buffer $\mathcal{B}_{F}^{h}$ stores transitions as
+[…]
+Here, $s_{T}$ is the terminal state, $\sum_{j}r_{j}$ is the cumulative reward collected until reaching the subgoal or waypoint, and $t_{\mathrm{wp}}$ is the time step at which $\mathrm{wp}_{\mathrm{final}}$ is reached.
+In FER, a success transition records the full cumulative return up to $t^{\prime}$ and continues the episode, a stop-on-failure transition assigns zero return and sets the next state to $s_{T}$, which immediately truncates the episode so all future returns are forfeited and unreachable or unreliable subgoals are discouraged, and a partial-success transition records $\mathrm{wp}_{\mathrm{final}}$ with its accumulated return, localizing how far the attempt progressed and crediting only the reliably executed portion.
+Together, these signals delineate the reachability frontier, identify unreliable subgoals, and provide consistent high-level targets that reduce unnecessary high-level decisions.
+This setup separates reliable from unreliable subgoals and suppresses failure-inducing actions, but it can induce conservatism and reduce exploratory coverage because the agent learns to avoid regions associated with failure.
+To counterbalance this effect, we introduce a decoupled high-level controller comprising a high-level policy $\pi^{h}$ for exploitation trained on $\mathcal{B}_{F}^{h}$ and a complementary exploration policy $\pi^{\mathrm{exp}}$ that promotes coverage.
+As in prior graph-based HRL, we assume the goal space $\mathcal{G}$ is known for subgoal selection of both $\pi^{h}$ and $\pi^{\mathrm{exp}}$, and SSE introduces no additional assumptions.
+Detailed formulations follow.
+High-level Policy $\pi^{h}$: In prior work, subgoals are typically selected with Gaussian policies with small noise.
+Although adequate for those methods, this concentrates exploration near the current maximum subgoal.
+As noted above, we aim for a high-level policy that can target any point in the goal space, thereby broadening exploration.
+We therefore define an $\epsilon$-greedy $\pi^{h}$ as
+$\pi^{h}(\tilde{g}_{t}\mid s_{t},g)=\begin{cases}\tilde{g}_{\max,t}:=\arg\max_{\tilde{g}\in\mathcal{G}}Q^{h}(s_{t},\tilde{g},g)&\text{with probability }1-\epsilon,\\ \tilde{g}_{\mathrm{rand}}\sim\mathrm{Uniform}(\mathcal{G})&\text{with probability }\epsilon,\end{cases}$ (3)
+where $\tilde{g}_{\max,t}$ is the greedy subgoal, in practice generated by the actor network trained to choose the maximum of $Q^{h}$, $\tilde{g}_{\mathrm{rand}}$ is sampled uniformly from $\mathcal{G}$ to ensure persistent global exploration, and $Q^{h}$ is trained off-policy using $\mathcal{B}_{F}^{h}$.
+This formulation allows $\pi^{h}$ to select diverse points in the goal space independently of the agent’s current location, which is crucial for broader exploration.
+Exploration Policy $\pi^{\mathrm{exp}}$: To promote exploration, $\pi^{\mathrm{exp}}$ targets low-density, underexplored regions of the goal space.
+To do this, we first identify novel regions $V_{\mathrm{novel}}\subset\mathcal{G}$, for which we employ either a grid-based estimator or a model-based approach.
+In this paper, we consider both: the grid-based estimator is used for simplicity in 2D/3D goal spaces that are widely adopted in real-world settings (e.g., position goal spaces), whereas the model-based approach, although more complex, is adopted to facilitate extension to higher-dimensional goal spaces.
+The exploration policy $\pi^{\mathrm{exp}}$ is defined as:
+$\pi^{\mathrm{exp}}(\tilde{g}_{t}\mid s_{t},g)=\begin{cases}g&\text{with probability }\tfrac{1}{3},\\ \tilde{g}_{\max,t}&\text{with probability }\tfrac{1}{3},\\ \tilde{g}_{\mathrm{novel}}\sim\mathrm{Uniform}(V_{\mathrm{novel}})&\text{with probability }\tfrac{1}{3}.\end{cases}$ (4)
+where the novel region $V_{\mathrm{novel}}$ is defined in one of two ways, depending on whether we use the grid-based or the model-based learning scheme, that is, $V_{\mathrm{novel}}=V_{\mathrm{novel}}^{\mathrm{grid}}$ or $V_{\mathrm{novel}}=V_{\mathrm{novel}}^{\mathrm{model}}$.
+The grid-based novel region is defined as $V_{\mathrm{novel}}^{\mathrm{grid}}:=C_{\mathcal{G}}^{m_{\mathrm{novel}}}$, where $C_{\mathcal{G}}^{m}$ denotes the grid cell of size $d_{\mathcal{G}}$ that partitions the goal space $\mathcal{G}$, $m$ indexes the grid cells, and $m_{\mathrm{novel}}=\arg\min_{m}N(C_{\mathcal{G}}^{m})$ is the index of the cell with the smallest visitation count $N(C_{\mathcal{G}}^{m})$.
+The model-based novel region is defined as $V_{\mathrm{novel}}^{\mathrm{model}}:=\left\{\arg\max_{v\in V}\lVert f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\rVert\right\},$ where $f_{\zeta}$ is the learner network with parameter $\zeta$, and $f_{\zeta_{\mathrm{targ}}}$ is the initialized target network proposed by (Burda et al., 2018) to describe novel regions based on prediction error.
+The learner $f_{\theta}$ is trained to minimize the prediction error for node samples.
+Then, the novel subgoal $\tilde{g}_{\mathrm{novel}}$ is uniformly sampled from $V_{\mathrm{novel}}$.
+Both $\pi^{h}$ and $\pi^{\mathrm{exp}}$ operate under the same SSE mechanism, where episodes continue only upon successful subgoal completion and terminate on failure.
+To control exploration, we sample from $\pi^{\mathrm{exp}}$ early in training and then gradually mix it with $\pi^{h}$ using a ratio $\eta:(1-\eta)$, where $\eta$ controls exploration strength.
+This balanced scheme preserves coverage, accelerates the discovery of reachable subgoals, and improves high-level generalization.
+Implementation details and parameters such as $\eta$ are described in  B.2
 Figure 2: Initial subgoals at $t=0$ selected by $\pi^{h}$ and $\pi^{\mathrm{exp}}$, with corresponding Ant agent trajectories at (a) early, (b) intermediate, and (c) final training stages in the U-maze task.
@@
 Over time, unreachable areas are excluded from subgoal candidates via SSE.
-While SSE encourages precise goal-directed behavior, its strict reward structure can limit exploration.
-To address this, we introduce a decoupled exploration policy $\pi^{\mathrm{exp}}$ that actively explores under-visited regions of the goal space.
-At each step $t$, $\pi^{\mathrm{exp}}$ selects a subgoal according to:
-$\pi^{\mathrm{exp}}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}g&\text{with probability }\frac{1}{3},\\ \tilde{g}_{\max,t}&\text{with probability }\frac{1}{3},\\ \text{randomly selected }\tilde{g}_{\mathrm{novel}}\in C_{\mathcal{G}}^{m_{\mathrm{novel}}}&\text{with probability }\frac{1}{3},\end{cases}$ (4)
-where $\tilde{g}_{\mathrm{novel}}$, referred to as the novel subgoal, is sampled from the least visited region $C_{\mathcal{G}}^{m_{\mathrm{novel}}}$ among $M$ grid cells $C_{\mathcal{G}}^{m}$ of size $d_{\mathcal{G}}$ that partition the goal space, $\mathcal{G}=\bigcup_{m=1}^{M}C_{\mathcal{G}}^{m}$.
-Here, $m_{\mathrm{novel}}$ is determined by $m_{\mathrm{novel}}=\arg\min_{m}N(C_{\mathcal{G}}^{m})$, where $N(C_{\mathcal{G}}^{m})$ denotes the number of episodes in which grid $C_{\mathcal{G}}^{m}$ has been visited.
-The policy $\pi^{\mathrm{exp}}$ follows the same SSE mechanism as $\pi^{h}$ and stores its transitions in the high-level buffer $\mathcal{B}^{h}$.
-This strategy promotes broad and effective exploration by encouraging visits to final goals, high-value subgoals, and novel subgoals in a balanced manner.
-Early in training, we rely solely on $\pi^{\mathrm{exp}}$ and gradually mix it with $\pi^{h}$ using a target ratio of $\eta:(1-\eta)$, where $\eta$ is the exploration control factor.
-By improving coverage of promising and under-explored regions without additional learning, $\pi^{\mathrm{exp}}$ enhances subgoal accessibility and high-level generalization.
 To illustrate the behavior of the proposed method, Fig.
 2 shows grid-wise visitation in the goal space and low-level trajectories toward various initial subgoals at three stages of training: (a) early (10K steps), (b) intermediate (150K steps), and (c) final (500K steps).
-The environment is a U-maze where a MuJoCo [43] Ant agent navigates to the final goal $g$ located at the upper left of the map.
-In (a), most of the goal space remains unexplored, and initial subgoals selected by $\pi^{\mathrm{exp}}$ and $\pi^{h}$, including the final goal, random subgoals, and novel regions, result in wide-ranging trajectories that promote broad exploration.
+The environment is a U-maze where a MuJoCo (Todorov et al., 2012) Ant agent navigates to the final goal $g$ located at the upper left of the map.
+In (a), most of the goal space remains unexplored, and initial subgoals selected by $\pi^{\mathrm{exp}}$ (grid-based) and $\pi^{h}$, including the final goal, random subgoals, and novel regions, result in wide-ranging trajectories that promote broad exploration.
 As training progresses in (b), the agent expands its coverage across the map, and in (c), it consistently reaches the goal $g$.
@@
 4.3 Failure-Aware Path Refinement
-In the SSE framework, reliable subgoal execution is crucial since each subgoal must be reached within a single high-level step.
-As described in Section 2, graph-based methods use Dijkstra’s algorithm to compute waypoint paths from $\phi(s_{t})$ to $\tilde{g}_{t}$ on a graph $G=(V,E)$, with edge distances defined by $d$ in Eq.
-(1).
-However, this distance does not capture failure cases such as collisions or getting stuck, leading agents to often fail even along the shortest path.
-We observe that such failure-prone regions significantly hinder subgoal success.
+In the SSE framework, reliable subgoal execution is crucial as each subgoal must be reached within a single high-level step.
+As described in Section 2, graph-based methods use Dijkstra’s algorithm to compute waypoint paths from $\phi(s_{t})$ to $\tilde{g}_{t}$ on a graph $G=(V,E)$, with edge distances defined by $d$ in equation 1.
+However, this distance ignores failure cases like collisions or getting stuck, causing agents to fail even on the shortest path.
+We observe these failure-prone regions significantly hinder subgoal success.
 To address this, we introduce a failure-aware path refinement strategy that increases edge costs in unreliable regions, steering the planner toward safer alternatives.
-Using the grid cells $C_{\mathcal{G}}^{m}$, we define $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ as the number of episodes where the agent fails to reach a subgoal located outside the current cell, i.e., $\tilde{g}_{t}\in C_{\mathcal{G}}^{m^{\prime}}$ with $m^{\prime}\neq m$, but the episode terminates within $C_{\mathcal{G}}^{m}$.
-This indicates failure to exit the current region.
-To discourage such behavior, we increase traversal costs in these areas by defining the failure ratio as $\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})=N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})/N(C_{\mathcal{G}}^{m})$, and refine the edge distance from node $v_{1}$ to $v_{2}$ as:
-$\tilde{d}(v_{1}\rightarrow v_{2})=d(v_{1}\rightarrow v_{2})\times\max\left(1,c_{\mathrm{dist}}\cdot\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})\right),\quad\forall v_{2}\in C_{\mathcal{G}}^{m},\quad\forall m,$ (5)
-where $d$ is the original edge distance defined in Eq.
-(1), and $c_{\mathrm{dist}}>1$ is a scaling factor.
-A higher failure ratio increases the adjusted distance, encouraging Dijkstra’s algorithm to avoid unreliable regions.
-If no alternative exists, the original path is still taken to preserve subgoal reachability.
+For each edge from node $v_{1}$ to node $v_{2}$, we define the failure ratio $\mathrm{ratio}_{\mathrm{fail}}$ for the target node $v_{2}$, and refine the edge by increasing its distance as the failure ratio increases as
+$\tilde{d}(v_{1}\rightarrow v_{2})=d(v_{1}\rightarrow v_{2})\times\max\left(1,c_{\mathrm{dist}}\cdot\mathrm{ratio}_{\mathrm{fail}}(v_{2})\right),~\forall v_{1},~v_{2}\in V,$ (5)
+where $d$ is the original edge distance from equation 1, and $c_{\mathrm{dist}}>1$ is a scaling factor.
+For consistency with our exploration policy, the failure ratio for the target node is also defined in two ways, following the grid-based and model-based approaches: $\mathrm{ratio}_{\mathrm{fail}}=\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{grid}}$ or $\mathrm{ratio}_{\mathrm{fail}}=\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{model}}$.
+The grid-based failure ratio is defined as $\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{grid}}(v_{2}):=N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})/N(C_{\mathcal{G}}^{m}),$ where $v_{2}\in C_{\mathcal{G}}^{m}$, $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ is the number of failure episodes in the cell $C_{\mathcal{G}}^{m}$.
+The model-based failure ratio is defined as $\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{model}}(v_{2}):=F_{\xi}(v_{2}),$ where $F_{\xi}$ is a failure predictor network with parameters $\xi$ trained over node samples using a cross-entropy objective, assigning label $1$ to nodes from failed trajectories and label 0 to nodes from successful trajectories.
+To ensure sufficient low-level policy competence, the failure-aware path refinement is activated only after $\lambda_{\mathrm{count}}$ successful visits.
+A higher failure ratio increases the adjusted edge distance, encouraging Dijkstra’s algorithm to avoid unreliable regions.
+As described above, we instantiate SSE with two variants that differ in their exploration policy and failure-aware path refinement, namely a grid-based approach and a model-based approach.
+We refer to SSE with the grid-based variant as SSE (Grid) and SSE with the model-based variant as SSE (Model), and we empirically compare these two variants to analyze how the choice of approach affects learning performance.
 Fig.
-3 illustrates the effect of the proposed path refinement in a bottleneck environment.
+3 illustrates the effect of the proposed grid-based path refinement in a bottleneck environment.
 In (a), without refinement, the agent repeatedly follows the shortest path through a narrow corridor near a wall, often resulting in failure.
@@
 4, with a condensed version of the algorithm provided in Algorithm 1.
-The full algorithm and implementation details, including the graph construction method and training procedures, are provided in Appendix B.
+The full algorithm and implementation details, including the graph construction and training losses for RL training and model-based approach are provided in Appendix B.
 Figure 3: Comparison of agent trajectories (blue lines) in a map with a bottleneck: (a) without path refinement and (b) with the proposed path refinement.
@@
     for each high-level selection step do
-                   Sample a subgoal $\tilde{g}_{t}$ from $\pi^{h}$ or $\pi^{\mathrm{exp}}$ with ratio $\eta:(1{-}\eta)$
-                  Plan the waypoint path from $\phi(s_{t})$ to $\tilde{g}_{t}$
-                  Roll out low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ $\forall i$
-                  if $\|\phi(s_{t+k_{t}})-\tilde{g}_{t}\|<\lambda$ (success) then
-                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})$
+         Sample a subgoal $\tilde{g}_{t}$ from $\pi^{h}$ and $\pi^{\mathrm{exp}}$
+        Plan a waypoint path from $\phi(s_{t})$ to $\tilde{g}_{t}$
+        Roll out low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ $\forall i$:
+       if $\|\phi(s_{t^{\prime}})-\tilde{g}_{t}\|<\lambda$ (success) then
+            Store the success transition and continue
        else
-                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})$ and terminate
+            Store the failure and partial-success transitions, then terminate the episode
          end if
-                  $N_{\mathrm{fail}}{(C_{\mathcal{G}}^{m})}+=1$ if $\phi(s_{t+k_{t}})\in C_{\mathcal{G}}^{m}$
       end for
-            $N(C_{\mathcal{G}}^{m})+=1$ for all visited grids $C_{\mathcal{G}}^{m}$
    end for
+  Count $N(C_{\mathcal{G}}^{m})$ and $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})~\forall m$ (Grid-based)
+  Update models $f_{\zeta}$ and $F_{\xi}$ (Model-based)
   Update $Q^{h}$, $Q^{l}$, $\pi^{h}$, $\pi^{l}$ via off-policy RL
@@
 5 Experiments
-In this section, we evaluate the proposed SSE framework on 9 challenging long-horizon tasks, including 5 AntMaze environments (U-maze, $\pi$-maze, AntMazeComplex, AntMazeBottleneck, and AntMazeDoubleBottleneck).
-Among these, U-maze has a simple layout, AntMazeBottleneck features narrow corridors, and the others present more complex structures.
+In this section, we evaluate our SSE framework on 9 challenging long-horizon tasks, including 5 AntMaze environments (U-maze, $\pi$-maze, AntMazeComplex, AntMazeBottleneck, and AntMazeDoubleBottleneck).
+These range from simple layouts (U-maze) to complex structures with narrow corridors (AntMazeBottleneck).
 We also assess 2 KeyChest tasks (AntKeyChest, AntDoubleKeyChest), where the agent must collect 1 or 2 keys before reaching the final goal, even though the keys are not explicitly defined as goals.
@@
 KeyChest and DoubleGoal tasks require intermediate objectives, making them ideal for testing high-level planning.
-Fig.
-5 provides visualizations, with further details presented in Appendix C.
+See Fig.
+5 for visualizations and Appendix C for details.
 While the main experiments focus on fixed-goal settings, additional comparisons for random-goal setups are included in Appendix D.1.
@@
 Figure 6: Performance comparison on various long-horizon environments
-We compare SSE with a range of baselines, including both hierarchical RL and recent graph-based methods.
-Specifically, we evaluate 2 HRL approaches: HIRO [40], which improves sample efficiency via hindsight goal relabeling, and HRAC [16], which penalizes subgoal selection based on reachability.
-We also include 3 graph-based HRL methods: HIGL [12], which applies intrinsic penalties over a graph, DHRL [19], which finds a path and gives waypoints to low-level as a goal via graph palnning, and NGTE [20], which expands graphs using novelty to better address fixed-goal settings.
-Additionally, we consider 2 graph-based methods without explicit high-level policies: PIG [42], which integrates graph structures into imitation learning to skip redundant subgoal actions, and BEAG [15], which uses breadth exploration with imaginary landmarks for goal-reaching.
-SSE is evaluated with the best hyperparameter settings ($c_{\mathrm{dist}}$, $d_{\mathcal{G}}$, $\eta$) from ablation studies, while all baselines use author-provided implementations.
+We compare SSE with a range of hierarchical RL and recent graph-based methods.
+Specifically, we evaluate 2 HRL approaches: HIRO (Nachum et al., 2018b), which improves sample efficiency via hindsight goal relabeling, and HRAC (Zhang et al., 2020), which penalizes subgoal selection based on reachability.
+We also include 3 graph-based HRL methods: HIGL (Kim et al., 2021), which applies intrinsic penalties over a graph, DHRL (Lee et al., 2022), which finds a path and gives waypoints to low-level as a goal via graph planning, and NGTE (Park et al., 2024), which expands graphs using novelty to better address fixed-goal settings.
+Additionally, we consider 2 graph-based methods without explicit high-level policies: PIG (Kim et al., 2023), which integrates graph structures into imitation learning to skip redundant subgoal actions, and BEAG (Yoon et al., 2024), which uses breadth exploration with imaginary landmarks for goal-reaching.
+For our proposed framework, we evaluate two variants: a grid-based implementation SSE (Grid), which offers a simple implementation in practical 2D/3D goal spaces, and a learning-based implementation SSE (Model), which is designed to be scalable to higher-dimensional goal spaces, as described in Section 4.
+Both variants are evaluated with the best hyperparameter settings ($c_{\mathrm{dist}}$, $d_{\mathcal{G}}$, $\eta$) from ablation studies, while all baselines use author-provided implementations.
+In FER, the threshold $\lambda$ used to determine subgoal reachability is the same as the threshold used to define the low-level reward in Section 2.2.
+We use the same $\lambda$ values as the baseline methods to ensure a fair and consistent comparison.
 Detailed descriptions of each algorithm, along with the hyperparameter configurations for our proposed method, are provided in Appendix C.
 Fig.
-6 shows success rate comparisons across all tasks, where solid lines indicate the mean over 5 seeds and shaded areas denote standard deviations.
-From the result, SSE consistently outperforms both graph-based and conventional HRL methods across all benchmarks.
+6 shows mean success rates over 5 seeds (solid lines) with standard deviations (shaded).
+From the results, both SSE variants consistently outperform other graph-based and conventional HRL methods across all benchmarks.
 Conventional HRL methods (DHRL, HIRO, PIG, HRAC, and HIGL) rely on random goal sampling for exploration but often struggle with fixed-goal tasks, while occasionally succeeding in random-goal setups provided in Appendix D.1, highlighting the increased challenge of fixed goals due to limited exploration opportunities.
-In relatively simple environments such as U-maze, $\pi$-maze, and AntMazeComplex, baseline methods like BEAG and NGTE demonstrate reasonable performance, but SSE typically converges more quickly.
+In relatively simple environments such as U-maze, $\pi$-maze, and AntMazeComplex, baseline methods like BEAG, grid-based nodes, and NGTE demonstrate reasonable performance, but SSE typically converges more quickly.
 In bottleneck environments, SSE further excels by using failure-aware path refinement to avoid unstable regions as shown in Fig.
@@
 In more complex tasks like KeyChest and ReacherWallDoubleGoal, which require reaching intermediate objectives, baseline methods largely fail.
-This is because conventional HRL suffers from long high-level horizons, and goal-centric methods without high-level decision-making, such as BEAG, cannot reason about intermediate targets.
+Conventional HRL suffers from long high-level horizons, and goal-centric methods without high-level decision-making, such as BEAG, cannot reason about intermediate targets.
 In contrast, SSE mitigates these issues by strictly enforcing subgoal completion, reducing high-level decision steps.
-These results highlight the robustness, efficiency, and generalization of the proposed SSE framework.
-For practical comparison, we also evaluate the computational complexity against major baselines in Appendix D.2.
-The results show that our method achieves lower complexity per iteration, demonstrating its superiority in terms of computational efficiency.
+These results highlight the versatility, efficiency, and generalization of the proposed SSE framework.
+While both variants succeed in all tasks, SSE (Grid) converges faster due to lower overhead.
+In contrast, SSE (Model) is designed for scalability to high-dimensional spaces where discretization is infeasible, despite higher training complexity.
+In addition, extended evaluations on AntPush, AntFall, and Ant4Rooms in Appendix E.3 confirm that SSE consistently matches or exceeds baselines.
+Furthermore, computational analysis in Appendix D.2 demonstrates that SSE achieves lower complexity per iteration than baselines.
 5.2 Further Analysis and Ablation Studies
+Figure 7: Trajectory analysis for SSE subgoals $\tilde{g}_{\max}$ in AntDoubleKeyChest at: (a) early stage, (b) reaches first key, (c) collects both keys, (d) reaches goal after collecting both keys (task success).
+In the following analysis and ablation studies, we focus on the grid-based approach and refer to SSE (Grid) simply as SSE, as it consistently achieves better performance and offers a richer set of hyperparameters for analysis and implementation.
 Fig.
@@
 As a result, it can solve complex multi-goal tasks using a minimal number of high-level steps, which highlights the effectiveness of SSE in long-horizon environments that require sequential decision-making for the high-level policy.
-Figure 7: Trajectory analysis for SSE subgoals $\tilde{g}_{\max}$ in AntDoubleKeyChest at: (a) early stage, (b) reaches first key, (c) collects both keys, (d) reaches goal after collecting both keys (task success).
+Notably, SSE solves this sequential task without an explicit curriculum.
+The reliability of its high-level policy, learned via FER, combined with an augmented state including key-possession flags, enables the agent to autonomously discover the required sequence of sub-objectives.
 (a) Component evaluation
@@
 Figure 8: Ablation study on AntDoubleKeyChest environment
-We conduct an ablation study on AntDoubleKeyChest to evaluate the contribution of each component in SSE and the impact of key hyperparameters, as shown in Fig.
+We conduct an ablation study on AntDoubleKeyChest to evaluate the contribution of each component in SSE and to analyze the impact of key hyperparameters, as shown in Fig.
 8.
-For the component analysis, we consider three variants: (1) SSE without path refinement (PR), which disables the failure-aware cost adjustment in graph planning; (2) SSE without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, $g$, or $\pi^{\mathrm{exp}}$, where we remove each subgoal type individually or disable the exploration policy altogether; and (3) SSE with a fixed low-level step (FLS), which replaces strict subgoal enforcement with a fixed-step high-level policy as in prior HRL methods.
-All variants lead to degraded performance, with the fixed-step setting failing to learn entirely, emphasizing the importance of strict subgoal execution and exploratory diversity.
-For hyperparameter sensitivity, setting $c_{\mathrm{dist}}=5$ achieves the optimal balance between penalizing failure-prone paths and avoiding excessive detours.
-Smaller values have little effect, while larger ones result in inefficient routing.
-Grid resolution remains stable, but overly fine grids (e.g., $d_{\mathcal{G}}=1$) slow learning due to increased planning costs.
-For the exploration ratio $\eta$, a value of 0.2 proves optimal, as lower values limit coverage too early and higher values delay the convergence of $\pi^{h}$.
-Despite some sensitivity, SSE consistently outperforms all baselines, demonstrating the SSE’s superiority.
-Additional trajectory analyses and ablation studies in other key environments are presented in Appendix E, further demonstrating the superiority of the proposed method.
+For the component analysis, we consider five variants: (1) SSE with FPS, which replaces the grid-based landmarks with Farthest Point Sampling Kim et al.
+(2021); (2) SSE with HER, which replaces strict subgoal execution with a conventional high-level policy that uses HER; (3) SSE without $\mathcal{B}_{F}^{h}$, which disables FER by using a standard replay buffer; (4) SSE without $\pi^{\mathrm{exp}}$, which disables the exploration policy; and (5) SSE without path refinement (PR), which disables the failure-aware adjustment.
+All variants exhibit degraded performance.
+Notably, the settings using HER or disabling FER fail to learn entirely, emphasizing the importance of strict subgoal execution and the consistent high-level signals provided by FER.
+SSE with FPS eventually succeeds, though less efficiently than the grid version, while still outperforming FPS-based baselines.
+This confirms that our framework remains effective regardless of the specific landmark selection strategy.
+Our hyperparameter analysis shows that setting $c_{\mathrm{dist}}=5$ achieves an optimal balance, though performance holds steady across a range of values.
+Similarly, performance also holds steady for grid resolution, though overly fine grids (e.g., $d_{\mathcal{G}}=1$) can slow learning.
+For the exploration ratio $\eta$, a value of 0.2 proves optimal.
+While hyperparameter tuning optimizes performance, SSE consistently outperforms all baselines.
+Additional analyses in other environments and sensitivity studies on the reachability threshold $\lambda$ and the high-level discount factor $\gamma^{h}$ are provided in Appendix E, further characterizing SSE.
 6 Limitation
-While SSE demonstrates strong performance across various long-horizon goal-conditioned tasks, it is currently limited to settings where the agent’s objective is solely to reach a specified goal.
-This assumption restricts its applicability to more complex tasks that require sequential or composite actions, such as robotic manipulation involving grasping and relocating objects, where multiple skill primitives may be necessary.
-Addressing such tasks would require extending the current framework to support a broader set of low-level skills and a high-level policy capable of selecting among them, presenting a promising direction for future work.
-Additionally, SSE relies on discretizing the goal space into grid cells to perform failure-aware path refinement.
-This approach is practical for environments with low-dimensional goal spaces such as 2D or 3D, commonly found in navigation and locomotion tasks.
-However, in high-dimensional goal spaces, the number of partitions increases rapidly, making exploration inefficient and coverage more difficult.
-To address this limitation, more advanced partitioning techniques such as clustering based on learned goal representations or adaptive sampling strategies could be considered to improve scalability in complex domains.
+Our framework introduces new hyperparameters, such as the exploration ratio $\eta$ and path refinement factor $c_{\mathrm{dist}}$, which require tuning.
+However, we find their effective ranges to be stable across diverse environments, and our ablation studies confirm that the framework maintains strong performance across a wide range of values, minimizing the overall tuning cost.
+While SSE also adds computational steps, its principle of early termination on subgoal failure yields significant efficiency gains.
+This prevents long, unproductive trajectories and results in a faster computation time compared to other recent methods, as quantitatively analyzed in Appendix D.2.
 7 Conclusion
-In this paper, We proposed Strict Subgoal Execution (SSE), a graph-based HRL framework designed to improve reliability and efficiency in long-horizon, goal-conditioned tasks.
-By enforcing single-step subgoal reachability, SSE enables robust high-level planning and significantly reduces decision horizons.
+In this paper, we proposed SSE, a graph-based HRL framework designed to improve reliability and efficiency in long-horizon, goal-conditioned tasks.
+By enforcing single-step subgoal reachability, SSE enables more direct and reliable high-level planning and significantly reduces decision horizons.
 The introduction of failure-aware path refinement and a decoupled exploration policy further enhances subgoal reliability and map coverage.
@@
 These results highlight the framework’s effectiveness in enabling stable and generalizable behavior, making it a promising approach for scalable hierarchical control in various long-horizon tasks.
-Appendix A Broader Impact
-The proposed SSE framework enhances the practicality and robustness of HRL in sparse-reward, long-horizon environments by eliminating fixed subgoal intervals and integrating failure-aware planning.
-This design reduces the dependency on reward shaping and broadens the applicability of HRL to real-world domains such as robotics, autonomous navigation, and long-term task planning.
-Furthermore, the graph-based structure of SSE improves interpretability, making it suitable for systems demanding transparency and reliability.
-However, its effectiveness still depends on careful design and validation; in safety-critical applications, poorly defined goals or sudden environmental shifts could lead to suboptimal behaviors.
-Nevertheless, SSE does not introduce specific societal or ethical concerns beyond standard RL considerations and poses no additional risks when applied responsibly.
+Acknowledgment
+This work was supported partly by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.
+RS-2022-II220469, Development of Core Technologies for Task-oriented Reinforcement Learning for Commercialization of Autonomous Drones), (No.
+RS-2025-25442824, AI Star-Fellowship Program (UNIST)), and (No.
+RS-2020-II201336, Artificial Intelligence Graduate School Support (UNIST)), and partly by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2025-23523191, LLM-Based Multi-Agent Reinforcement Learning for End-to-End Large Autonomous Swarm Control).
+Ethics statement
+This paper introduces a foundational algorithm, SSE, designed to improve the long-horizon planning capabilities of reinforcement learning agents.
+All experiments were conducted in standard, simulated robotics environments.
+As such, our research does not involve human subjects, sensitive or personally identifiable data, nor does it directly address systems that interact with people.
+Therefore, issues of data privacy, dataset bias, and fairness are not directly applicable to this work.
+Reproducibility statement
+We are committed to the reproducibility of our research.
+The complete source code for our proposed framework, SSE, and all experiments is included as an anonymized zip file in the supplementary materials.
+A detailed breakdown of the implementation, including network architectures and training procedures, can be found in Appendix B.
+All hyperparameters required to reproduce our results are provided in Appendix C.3, with common settings listed in Table 2 and environment-specific configurations in Table 3.
+Furthermore, Appendix C details the full experimental setup, including descriptions of the baseline algorithms, specifications for all environments (Table 1), and the hardware and software configurations on which the experiments were conducted.
+We believe these resources will enable the reproduction of our findings.
+Appendix A The Use of Large Language Models
+We utilized a large language model (LLM) as an assistive tool during the preparation of this manuscript.
+The LLM’s role was strictly limited to polishing the text, which includes improving clarity, conciseness, and correcting grammatical errors.
+The LLM was not used for research ideation.
+The authors have carefully reviewed and edited all content and take full responsibility for the scientific accuracy and integrity of this work.
+The LLM is not credited as an author.
 Appendix B Implementation Details
@@
 B.1 Graph Construction of SSE
-In this section, we describe how existing methods define the graph $G=(V,E)$ with the landmark node set $V$ and the edge set $E$ with edge distances $d$, and how the proposed SSE framework constructs this graph.
+In this section, we describe how existing methods define the graph $G=(V,E)$,which consists of a landmark node set $V$ and the edge set $E$ with edge distances $d$.
+We then explain how the proposed SSE framework constructs this graph.
 Landmark Node Set $V$ Construction
-Existing HRL methods such as DHRL [19] and NGTE [20] construct the landmark node set $V$ by selecting visited states during exploration.
+Existing HRL methods such as DHRL (Lee et al., 2022) and NGTE (Park et al., 2024) construct the landmark node set $V$ by selecting visited states during exploration.
 They employ the Farthest Point Sampling (FPS) algorithm to identify landmark nodes that are far apart from each other.
@@
 However, it is limited to visited states, meaning unexplored or infrequently visited areas cannot be selected as landmarks, slowing exploration in those regions.
-In contrast, BEAG [15] accelerates exploration by partitioning the goal space into a grid structure, creating landmark nodes at each grid intersection.
+In contrast, BEAG (Yoon et al., 2024) accelerates exploration by partitioning the goal space into a grid structure, creating landmark nodes at each grid intersection.
 This method allows for faster exploration by using virtual goal positions as landmarks, independent of visitation frequency.
 The structured grid layout enables more systematic and efficient exploration.
-To leverage this advantage, the proposed SSE framework adopts a grid-based strategy for landmark node selection inspired by BEAG.
+To leverage this advantage and ensure a structured, reproducible setup, SSE adopts a grid-based landmark selection strategy inspired by BEAG.
 Given a 2D goal space of size $x\times y$ and a grid size of $d_{\mathcal{G}}$, the landmark set $V$ is constructed as follows:
@@
 Definition of Edge Distance
-Given the landmark node set $V$, the edge set $E$ is defined as the collection of distances $d(v_{1}\rightarrow v_{2})$ between any two nodes $v_{1},v_{2}\in V$.
-For previous graph-based RL methods, the edge distance $d$ is computed as $d(v_{1}\rightarrow v_{2}):=\log_{\gamma^{l}}\left(1+(1-\gamma^{l})Q^{G}(v_{1},v_{2},\pi^{l})\right),$ where $\gamma^{l}$ is the low-level discount factor used for training the low-level policy $\pi^{l}$, and $Q^{G}$ is the value function estimating the traversal cost from $v_{1}$ to $v_{2}$, as described in Section 2.
-Existing HRL methods like DHRL [19] and NGTE [20] directly use the low-level value function $Q^{l}$, which is trained with a step-based reward of $-1$, to define the distance as the expected number of steps required for the low-level policy to navigate from $v_{1}$ to $v_{2}$.
-Although this method reflects actual navigation costs, it is sensitive to instability during $Q^{l}$ training, resulting in fluctuating edge distances.
-In contrast, BEAG [15] measures the distance using the Euclidean norm $d_{E}=||v_{1}-v_{2}||$, providing a stable but less accurate representation of traversal costs.
-To leverage the strengths of both approaches, SSE defines the edge distance $d(v_{1}\rightarrow v_{2})$ by combining $Q^{l}$ and $d_{E}$ as follows:
+[…]
 $d(v_{1}\rightarrow v_{2})=\frac{1}{2}\left[\log_{\gamma^{l}}\left(1+(1-\gamma^{l})Q^{l}(v_{1},v_{2},\pi^{l})\right)+\log_{\gamma^{l}}\left(1+(1-\gamma^{l})d_{E}(v_{1},v_{2})\right)\right].$ (7)
@@
 B.2 Detailed Loss Functions and Implementation of the SSE Framework
+RL training losses
 As described in Section 4, the proposed SSE framework is an HRL structure that employs a high-level policy $\pi^{h}$, an exploration policy $\pi^{\textrm{exp}}$, and a low-level policy $\pi^{l}$.
-The exploration policy does not require separate parameterization for high-level actions, while $\pi^{h}$ and $\pi^{l}$ are parameterized by $\theta^{h}$ and $\theta^{l}$, respectively, and represented as $\pi^{h}_{\theta^{h}}$ and $\pi^{l}_{\theta^{l}}$.
+The exploration policy does not require separate parameterization for high-level actions, whereas $\pi^{h}$ and $\pi^{l}$ are parameterized by $\theta^{h}$ and $\theta^{l}$, respectively, and are represented as $\pi^{h}_{\theta^{h}}$ and $\pi^{l}_{\theta^{l}}$.
 To evaluate these policies, SSE defines the parameterized high-level value function $Q^{h}_{\psi^{h}}$ and the low-level value function $Q^{l}_{\psi^{l}}$, where $\psi^{h}$ and $\psi^{l}$ are the respective parameters.
 As mentioned in Section 2, the high-level policy and value function are trained to maximize external rewards, while the low-level policy and value function are optimized to reach designated waypoints incrementally.
-To facilitate this, the low-level policy receives the following reward $r^{l}_{t}$ at each time step:
-$r^{l}_{t}=\begin{cases}0&\text{if }\|\phi(s_{t+1})-\mathrm{wp}_{i}\|<0.5,\textrm{ (agent reaches the current waypoint)}\\ -1&\text{otherwise},\end{cases}$ (8)
-where $\mathrm{wp}_{i}$ is the target waypoint at the current timestep $t$, and $(\mathrm{wp}_{1},\cdots,\mathrm{wp}_{n})$ represent the shortest path from $\phi(s_{t})$ to the subgoal $\tilde{g}_{t}$.
-The high-level and low-level policies, along with their value functions, are trained using the transitions stored in the high-level buffer $\mathcal{B}^{h}$ and the low-level buffer $\mathcal{B}^{l}$ through the TD3 algorithm [44], a standard off-policy RL method.
+To facilitate this, the low-level policy receives the reward $r^{l}_{t}$ at each time step, defined as follows:
+$r^{l}_{t}=\begin{cases}0&\text{if }\|\phi(s_{t+1})-\mathrm{wp}_{i}\|<\lambda,\textrm{ (agent reaches the current waypoint)}\\ -1&\text{otherwise},\end{cases}$ (8)
+where $\lambda$ is the reachability check threshold, $\mathrm{wp}_{i}$ is the target waypoint at the current timestep $t$, and $(\mathrm{wp}_{1},\cdots,\mathrm{wp}_{n})$ represent the shortest path from $\phi(s_{t})$ to the subgoal $\tilde{g}_{t}$.
+The high-level and low-level policies, along with their value functions, are trained using the transitions stored in the FER $\mathcal{B}_{F}^{h}$ and the low-level buffer $\mathcal{B}^{l}$ through the TD3 algorithm (Fujimoto et al., 2018), a standard off-policy RL method.
 The value function losses for high-level and low-level policies are defined as follows:
-$\small\mathcal{L}_{Q^{h}}(\psi^{h})$ $=\mathbb{E}_{B^{h}}\left[\left(Q^{h}(\tilde{s}_{t},\tilde{g}_{t})-\left(r^{h}_{t}+\gamma^{h}\min_{i=1,2}Q^{h}_{\bar{\psi}^{h}_{i}}(\tilde{s}_{t+k_{t}},\pi^{h}_{\theta^{h}}(\tilde{s}_{t+k_{t}},g))\right)\right)^{2}\right]$
+$\mathcal{L}_{Q^{h}}(\psi^{h})$ $=\mathbb{E}_{B_{F}^{h}}\left[\left(Q^{h}(s_{t},\tilde{g}_{t})-\left(r^{h}_{t}+\gamma^{h}\min_{i=1,2}Q^{h}_{\bar{\psi}^{h}_{i}}(s_{t^{\prime}},\pi^{h}_{\theta^{h}}(s_{t^{\prime}},g))\right)\right)^{2}\right]$
 $\mathcal{L}_{Q^{l}}(\psi^{l})$ $=\mathbb{E}_{B^{l}}\left[\left(Q^{l}(s_{t},a_{t})-\left(r^{l}_{t}+\gamma^{l}\min_{i=1,2}Q^{l}_{\bar{\psi}^{l}_{i}}(s_{t+1},\pi^{l}_{\theta^{l}}(s_{t+1},\mathrm{wp}_{i}))\right)\right)^{2}\right],$ (9)
-where $r^{h}_{t}=\sum_{j=t}^{t+k_{t}-1}r_{j}$ for a successful trajectory and $r^{h}_{t}=0$ for a failed trajectory, $\bar{\psi}^{h}$ and $\bar{\psi}^{l}$ are target network parameters updated via exponential moving average (EMA), and $\gamma^{h}$ and $\gamma^{l}$ are the discount factors for training the high-level and low-level policies, respectively.
+where $t^{\prime}$ denotes the termination time of the low-level path execution, which is variable as the step concludes only upon success or failure.
+In the case of a partial success, $t_{\mathrm{wp}}$ (used to store transitions in FER) indicates the time step when the agent reached the last successful waypoint $wp_{\mathrm{final}}$.
+The high-level reward is then defined as $r^{h}_{t}=\sum_{j=t}^{t^{\prime}-1}r_{j}$ for a successful trajectory and $r^{h}_{t}=0$ for a failed one.
+The terms $\bar{\psi}^{h}$ and $\bar{\psi}^{l}$ are target network parameters updated via exponential moving average (EMA), and $\gamma^{h}$ and $\gamma^{l}$ are the discount factors for training the high-level and low-level policies, respectively.
 The actor losses for optimizing the policies are defined as follows:
-$\mathcal{L}_{\pi^{h}}(\theta^{h})=-\mathbb{E}_{B^{h}}\left[Q^{h}_{\psi^{h}}(\tilde{s}_{t},\pi_{\theta^{h}}^{h}(\tilde{s}_{t},g))\right],\quad\mathcal{L}_{\pi^{l}}(\theta^{l})=-\mathbb{E}_{B^{l}}\left[Q^{l}_{\psi^{l}}(s_{t},\pi_{\theta^{l}}^{l}(s_{t},\mathrm{wp}_{i}))\right].$ (10)
-The parameters are optimized using the Adam optimizer [45] to minimize the respective loss functions.
+$\mathcal{L}_{\pi^{h}}(\theta^{h})=-\mathbb{E}_{B_{F}^{h}}\left[Q^{h}_{\psi^{h}}(s_{t},\pi_{\theta^{h}}^{h}(s_{t},g))\right],\quad\mathcal{L}_{\pi^{l}}(\theta^{l})=-\mathbb{E}_{B^{l}}\left[Q^{l}_{\psi^{l}}(s_{t},\pi_{\theta^{l}}^{l}(s_{t},\mathrm{wp}_{i}))\right].$ (10)
+The parameters are optimized using the Adam optimizer (Kingma, 2014) to minimize the respective loss functions.
 SSE distinguishes between the high-level discount factor $\gamma^{h}$ and the low-level discount factor $\gamma^{l}$, setting $\gamma^{l}=0.99$ as typical in RL, and $\gamma^{h}=0.4$ to limit return propagation across high-level steps, encouraging shorter path optimization.
 In the initial stages of training, rewards are sparse.
-To address this, the high-level buffer $\mathcal{B}^{h}$ is divided evenly, with half storing successful trajectories and the other half storing trajectories with zero reward.
+To address this, the FER $\mathcal{B}_{F}^{h}$ is divided evenly, with half storing successful trajectories and the other half storing trajectories with zero reward.
 This design improves learning signals from successful experiences.
-The exploration ratio, which controls the balance between the exploration policy and the high-level policy, starts at $1{:}0$ and decays by 0.05 per iteration until it reaches $\eta{:}(1{-}\eta)$, as outlined in Section 4.
+The exploration ratio, which controls the balance between the exploration policy and the high-level policy, starts at 1:01{:}0 and decays by 0.05 per iteration until it reaches $\eta{:}(1{-}\eta)$, as outlined in Section 4.
 This scheduling promotes exploration initially and shifts the focus to high-level learning as training progresses.
@@
 The complete SSE algorithm is provided in Algorithm 2.
+Model training losses for SSE (Model)
+We now describe the implementation details of the prediction network $f_{\zeta}$, which computes the prediction error used to determine the novel region in the model-based approach, and the failure predictor network $F_{\xi}$, which estimates the failure ratio of target nodes.
+In the model-based variant of the exploration policy, we estimate the novel region $V_{\mathrm{novel}}^{\mathrm{model}}$ using Random Network Distillation (RND) (Burda et al., 2018), which was originally proposed in standard RL to identify novel states based on prediction error.
+Following the RND formulation, we use a fixed target network $f_{\zeta_{\mathrm{targ}}}:V\rightarrow\mathbb{R}^{k}$ and a predictor network $f_{\zeta}:V\rightarrow\mathbb{R}^{k}$ with parameters $\zeta$.
+The predictor $f_{\zeta}$ is trained on visited sample nodes stored in the buffer $\mathcal{D}$ to regress the target output via the mean squared error objective
+$L_{\mathrm{RND}}(\zeta)=\mathbb{E}_{v\sim\mathcal{D}_{\mathrm{visited}}}\left[\left\|f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\right\|_{2}^{2}\right].$ (11)
+As the predictor network is trained to minimize the prediction error on visited nodes, the prediction error becomes small in frequently explored regions and remains large for underexplored nodes.
+We therefore use the squared prediction error $\left\|f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\right\|_{2}^{2}$ as a novelty score, and nodes with larger error are treated as more novel.
+By selecting nodes with high prediction error as exploration targets, the model-based variant is encouraged to visit underexplored regions of the goal space.
+For failure-aware refinement, we train a failure predictor network $F_{\xi}:V\rightarrow[0,1]$ parameterized by $\xi$, which outputs the estimated failure probability of each node.
+The network is trained using node samples stored in buffer $\mathcal{D}_{v}$, where each node $v$ is paired with a binary label $y\in\{0,1\}$ indicating whether the trajectory containing $v$ failed $(y=1)$ or succeeded $(y=0)$.
+We optimize $F_{\xi}$ with the binary cross-entropy loss
+$L_{F}(\xi)=-\mathbb{E}_{(v,y)\sim\mathcal{D}_{v}}\left[y\log(F_{\xi}(v))+(1-y)\log(1-F_{\xi}(v))\right].$ (12)
+Through this training, $F_{\xi}(v)$ approximates the empirical failure ratio at node $v$, which we use as the model-based failure estimate in our refinement scheme.
 Input: Graph $G=(V,E)$, goal $g$, mapping function $\phi$, threshold $\lambda$, exploration ratio $\eta$
-Initialize: Policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, buffers $B^{h}$, $B^{l}$, grid cells $C_{\mathcal{G}}$
+Initialize: Policies $\pi^{h}$, $\pi^{\mathrm{exp}}$, $\pi^{l}$, buffers $B_{F}^{h}$, $B^{l}$, grid cells $C_{\mathcal{G}}$ (Grid) or Models $f_{\zeta},F_{\xi}$ (Model)
 for each iteration do
@@
     for each high-level selection step do
-                   Sample a subgoal $\tilde{g}_{t}\sim\pi(\tilde{s}_{t},g)$
+         Sample a subgoal $\tilde{g}_{t}\sim\pi(s_{t},g)$
         Plan the waypoint path $\mathrm{wp}_{1:n}$ from $\phi(s_{t})$ to $\tilde{g}_{t}$ using Dijkstra’s algorithm over $G$ with $\tilde{d}$
        for $i=1,\cdots,n$ do
-                         Roll out the low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ to reach the waypoint $\mathrm{wp}_{i}$ within $k_{t}$ steps
-                        Store the $k_{t}$ transitions $(s_{t},\mathrm{wp}_{i},a_{t},r_{t},s_{t+1})$ into the low-level buffer $\mathcal{B}^{l}$
-                        Compute the reward sum: $r_{\mathrm{sum}}=\sum_{j=t}^{t+k_{t}-1}r_{j}$
+            Roll out the low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ to reach the waypoint $\mathrm{wp}_{i}$
+           Store the $t^{\prime}-t$ transitions $(s_{t},\mathrm{wp}_{i},a_{t},r_{t},s_{t+1})$ into the low-level buffer $\mathcal{B}^{l}$
+           Compute the reward sum: $r_{t}^{h}=\sum_{j=t}^{t^{\prime}-1}r_{j}$
+          Construct the FER:
           if $\|\phi(s_{t+1})-\tilde{g}_{t}\|<\lambda$ then
-                               Success: Store the high-level transition $(\tilde{s},g,\tilde{g}_{t},r_{\mathrm{sum}},\tilde{s}_{t+k_{t}})$ into $B^{h}$
+             Success: Store the success transition $(s,g,\tilde{g}_{t},r_{\mathrm{sum}},s_{t^{\prime}})$ into FER $B_{F}^{h}$
           else
-                               Failure: Store $(\tilde{s},g,\tilde{g}_{t},0,\tilde{s}_{T})$ into $B^{h}$
-                              $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})+=1$ for $m$ s.t. $\phi(s_{t+k_{t}})\in C_{\mathcal{G}}^{m}$
+             Failure: Store the stop-on-failure transition $(s,g,\tilde{g}_{t},0,s_{T})$ into FER $B_{F}^{h}$
+              Store the partial success transition $(s,g,\mathrm{wp}_{\mathrm{final}},\sum_{j=t}^{t_{\mathrm{wp}}-1}r_{j},s_{t_{\mathrm{wp}}})$ into FER $B_{F}^{h}$
               Terminate the episode
@@
       end for
-            $N(C_{\mathcal{G}}^{m})+=1$ for all visited cells $C_{\mathcal{G}}^{m}$
    end for
-      Update $\psi^{h}$, $\psi^{l}$, $\theta^{h}$, $\theta^{l}$ to minimize RL losses $\mathcal{L}_{Q^{h}}$, $\mathcal{L}_{Q^{l}}$, $\mathcal{L}_{\pi^{h}}$, $\mathcal{L}_{\pi^{l}}$, respectively
+  Update total visitation count $N(C_{\mathcal{G}}^{m})$ and failure count $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ for all $C_{\mathcal{G}}^{m}$ (Grid-based)
+  Update model parameters $\zeta$, $\xi$ to minimize $L_{\mathrm{RND}}(\zeta)$ and $L_{\mathrm{F}}(\xi)$ (Model-based)
+  Update $\psi^{h},\theta^{h}$ using samples from FER $B_{F}^{h}$ to minimize $\mathcal{L}_{Q^{h}}(\psi^{h}),\mathcal{L}_{\pi^{h}}(\theta^{h})$
+  Update $\psi^{l},\theta^{l}$ using samples from $\mathcal{B}^{l}$ to minimize $\mathcal{L}_{Q^{l}}(\psi^{l}),\mathcal{L}_{\pi^{l}}(\theta^{l})$
 end for
@@
 Appendix C.1 provides descriptions of the baseline algorithms along with links to their official code repositories.
-Appendix C.2 provides specifications of the nine environments shown in Fig.9, including their action and observation spaces, goal configurations, and episode horizons.
+Appendix C.2 provides the specifications for the nine environments shown in Fig.9, including their action and observation spaces, goal configurations, and episode horizons.
 The SSE-specific hyperparameter configurations for each environment are summarized in AppendixC.3.
 C.1 Details of Other Baselines
-  • HIRO [40] introduces a hierarchical architecture with relabeling of high-level transitions to account for changing low-level policies, thereby improving off-policy sample efficiency and stability.
+  • HIRO (Nachum et al., 2018b) introduces a hierarchical architecture with relabeling of high-level transitions to account for changing low-level policies, thereby improving off-policy sample efficiency and stability.
 Open-source code of HIRO is available at https://github.com/watakandai/hiro_pytorch
-  • HRAC [16] adds a learned adjacency constraint to ensure subgoal feasibility.
+  • HRAC (Zhang et al., 2020) adds a learned adjacency constraint to ensure subgoal feasibility.
 It penalizes high-level selections that attempt transitions deemed unreachable within a limited horizon, thereby guiding the agent to learn feasible subgoal structures.
 Open-source code of HRAC is available at https://github.com/trzhang0116/HRAC
-  • HIGL [12] incorporates a coverage-driven and novelty-driven landmark selection strategy.
+  • HIGL (Kim et al., 2021) incorporates a coverage-driven and novelty-driven landmark selection strategy.
 It performs graph-based planning via shortest paths and uses adjacency rewards to guide learning toward under-explored regions.
 Open-source code of HIGL is available at https://github.com/junsu-kim97/HIGL
-  • DHRL [19] constructs a goal graph using Farthest Point Sampling and learns high-level behavior by planning over the graph.
+  • DHRL (Lee et al., 2022) constructs a goal graph using Farthest Point Sampling and learns high-level behavior by planning over the graph.
 It emphasizes temporal abstraction and long-horizon planning via graph traversal and Q-learning.
 Open-source code of DHRL is available at https://github.com/jayLEE0301/dhrl_official
-  • BEAG [15] employs value-function-driven imaginary landmarks to facilitate exploration of unvisited areas.
+  • BEAG (Yoon et al., 2024) employs value-function-driven imaginary landmarks to facilitate exploration of unvisited areas.
 It estimates landmark distances from a learned value function without relying solely on previously visited states, enabling efficient generalization.
 Open-source code of BEAG is available at https://github.com/ml-postech/BEAG
-  • NGTE [20] drives novelty-based exploration by identifying frontier nodes (outposts) and prioritizing expansion toward less-visited regions of the goal graph, encouraging broad and diverse exploration.
+  • NGTE (Park et al., 2024) drives novelty-based exploration by identifying frontier nodes (outposts) and prioritizing expansion toward less-visited regions of the goal graph, encouraging broad and diverse exploration.
 Open-source code of NGTE is available at https://github.com/ihatebroccoli/NGTE
 C.2 Environmental Details
-We follow standard benchmarks and configurations widely adopted in prior hierarchical reinforcement learning studies [19, 20, 15], and introduce several new environments designed to evaluate high-level decision-making capabilities such as AntKeyChest or AntDoubleKeyChest.
+We follow standard benchmarks and configurations widely adopted in prior hierarchical reinforcement learning studies (Lee et al., 2022; Park et al., 2024; Yoon et al., 2024), and introduce several new environments designed to evaluate high-level decision-making capabilities such as AntKeyChest or AntDoubleKeyChest.
 The environments used in our experiments are visualized in Fig.
@@
 Action: 8-Dof
-[-4,20]$\times$
+[…]
 [-4,20]
-$1$ if goal reached
+11 if goal reached
 else 0
@@
 Action: 8-Dof
-[-4,36]$\times$
+[…]
 [-4,36]
-$1$ if goal reached
+11 if goal reached
 else 0
@@
 Action: 8-Dof
-[-4,52]$\times$
+[…]
 [-4,52]
-$1$ if goal reached
+11 if goal reached
 else 0
@@
 Action: 8-Dof
-[-4,20]$\times$
+[…]
 [-4,20]
-$1$ if goal reached
+11 if goal reached
 else 0
@@
 Action: 8-Dof
-[-4,20]$\times$
+[…]
 [-4,36]
-$1$ if goal reached
+11 if goal reached
 else 0
@@
 Action: 8-Dof
-[-4,36]$\times$
+[…]
 [-4,36]
-$1$ if key reached
-$5$ if goal reached
+11 if key reached
+55 if goal reached
 with key
@@
 Action: 8-Dof
-[-4,36]$\times$
+[…]
 [-4,36]
-$1$ if key1 reached
-$1$ if key2 reached
+11 if key1 reached
+11 if key2 reached
 with key1
-$5$ if goal reached
+55 if goal reached
 with two keys
@@
 Action: 7-Dof
-[-1,1]$\times$
-[-1,1]$\times$
+[…]
+[…]
 [-1,1]
 1 if goal reached Start:
-(0.99, --0.19, 0)
+(0.99, -0.19, 0)
 Goal:
-(0.6, 0.6, --0.1)
+(0.6, 0.6, -0.1)
 100
@@
 Action: 7-Dof
-[-1,1]$\times$
-[-1,1]$\times$
+[…]
+[…]
 [-1,1]
@@
 Start:
-(0.99, --0.19, 0)
+(0.99, -0.19, 0)
 Goal1:
-(0.4, 0.4, --0.1)
+(0.4, 0.4, -0.1)
 Goal2:
-(0.4, --0.8, --0.1)
+(0.4, -0.8, -0.1)
 200
@@
 Learning rates for each policy level, as well as the discount factors, were determined through a structured parameter search.
-Table 3 presents environment-specific hyperparameters, including the minimum epsilon threshold $\epsilon_{\text{min}}$, the path refinement scaling factor $c_{\mathrm{dist}}$, the subgoal success threshold $\lambda$, and the grid resolution $d_{\mathcal{G}}$.
+Table 3 presents environment-specific hyperparameters, including the minimum epsilon threshold $\epsilon_{\text{min}}$, the path refinement scaling factor $c_{\mathrm{dist}}$, the reachability check threshold $\lambda$, and the grid resolution $d_{\mathcal{G}}$.
+As mentioned in Section 2.2, the threshold $\lambda$ is the standard, environment-provided value used to check for low-level reachability (see equation 8), which is also used consistently across baselines.
 These values were selected based on targeted parameter searches conducted for each environment.
@@
 Table 3: Scenario-specific hyperparameters for SSE.
-Scenario $\boldsymbol{\epsilon_{\text{min}}}$ $\boldsymbol{c_{\mathrm{dist}}}$ $\boldsymbol{\lambda}$ $\boldsymbol{d_{\mathcal{G}}}$ $\eta$
+Scenario $\bm{\epsilon_{\text{min}}}$ $\bm{c_{\mathrm{dist}}}$ $\bm{\lambda}$ $\bm{d_{\mathcal{G}}}$ $\eta$
 AntMaze
-U-maze 0.1 5.0 2.0 2 0.1
-AntMazeBottleneck 0.1 5.0 2.0 2 0.1
-$\pi$-maze 0.1 5.0 2.0 2 0.2
-AntMazeComplex 0.1 5.0 2.0 2 0.2
-AntMazeDoubleBottleneck 0.1 10.0 2.0 2 0.1
-AntKeyChest 0.1 5.0 2.0 2 0.2
-AntDoubleKeyChest 0.2 5.0 2.0 2 0.2
+U-maze 0.1 5.0 0.5 2 0.1
+AntMazeBottleneck 0.1 5.0 0.5 2 0.1
+$\pi$-maze 0.1 5.0 0.5 2 0.2
+AntMazeComplex 0.1 5.0 0.5 2 0.2
+AntMazeDoubleBottleneck 0.1 10.0 0.5 2 0.1
+AntKeyChest 0.1 5.0 0.5 2 0.2
+AntDoubleKeyChest 0.2 5.0 0.5 2 0.2
 Reacher
-ReacherWall 0.1 5.0 0.15 1 0.2
-ReacherWallDoubleGoal 0.1 5.0 0.15 1 0.2
+ReacherWall 0.1 5.0 0.25 1 0.2
+ReacherWallDoubleGoal 0.1 5.0 0.25 1 0.2
 Appendix D Additional Comparative Experiments
@@
 D.1 Performance Comparison under Random Goal Setups
+Figure 10 presents experimental results under a random goal setting where goals are sampled uniformly from the entire valid state space during training.
+This setup contrasts with the fixed-goal scenarios discussed in the main text.
+The reward signal remains sparse and poses a significant challenge for goal discovery and policy optimization.
+Graph-based methods such as BEAG, NGTE, and SSE maintain strong performance even under random goal sampling.
+These methods autonomously expand their subgoal graph during exploration and achieve success rates comparable to those in the fixed-goal setting.
+DHRL also benefits from the randomized goal distribution, particularly in simpler environments like U-maze and AntMazeBottleneck where the agent encounters training signals more frequently.
+Conversely, methods such as HIRO, HIGL, and HRAC exhibit limited progress in most environments under random-goal conditions.
+This performance drop stems from their reliance on fixed-frequency subgoal selection and the lack of structured exploration mechanisms suitable for sparse reward settings.
+In environments requiring multi-stage reasoning such as AntKeyChest, SSE is the only method that consistently discovers the key and solves the full task.
+While NGTE is hierarchical and exploration-driven, it only occasionally learns to acquire the key and exhibits a very low overall success rate under this configuration.
 Figure 10: Performance comparison in random goal setting
-In Fig.
-10, we compare experimental results under a random goal setting, where the goal is sampled uniformly from the entire valid state space during training, unlike the fixed-goal scenarios discussed in the main text.
-The reward signal remains sparse, posing a significant challenge for goal discovery and policy optimization.
-Graph-based methods such as BEAG, NGTE, and SSE, which can autonomously expand their subgoal graph during exploration, maintain strong performance even under random goal sampling, achieving success rates comparable to those in the fixed-goal setting.
-DHRL also benefits from the randomized goal distribution, particularly in simpler environments like U-maze and AntMazeBottleneck, where the training signal is more frequently encountered.
-On the other hand, methods such as HIRO, HIGL, and HRAC exhibit limited progress in most environments under random-goal conditions, due to their reliance on fixed-frequency subgoal selection, sparse reward setting, and lack of structured exploration mechanisms.
-In environments requiring multi-stage reasoning such as AntKeyChest, SSE is the only method that consistently discovers the key and solves the full task.
-NGTE, while hierarchical and exploration-driven, occasionally learns to acquire the key, but overall exhibits a very low success rate under this random goal configuration.
 D.2 Computational Complexity Comparison
-Table 4 reports the average per-episode computation time and the number of episodes required to reach a 60% success rate with NGTE and BEAG.
-We compare against NGTE and BEAG because they are the only baselines that succeed in at least one of the considered fixed-goal settings, making them the most relevant references for evaluating performance and sample efficiency in sparse reward environments.
-SSE achieves faster convergence across tasks due to its early termination of episodes upon subgoal failure and a reduced number of transitions per episode, which minimizes unnecessary computation and accelerates learning.
-This advantage is particularly evident in long-horizon tasks such as AntKeyChest, where conventional HRL methods consume many timesteps even during failed attempts and require extended horizons for high-level planning.
-In contrast, in shorter-horizon scenarios such as AntMazeBottleneck, the computational benefits of SSE are less significant since subgoal transitions and failure terminations occur less frequently.
-The operations introduced by SSE, including the construction of $\tilde{s}$, evaluation of subgoal completion, and uniform sampling, are lightweight and have linear time complexity.
-As a result, SSE imposes minimal computational overhead while maintaining stable training dynamics.
+Table 4 reports the average per-episode computation time and the number of episodes required to reach a 60% success rate.
+We compare SSE against NGTE and BEAG as they are the only baselines that succeed in at least one of the considered fixed-goal settings.
+This makes them the most relevant references for evaluating performance and sample efficiency in sparse reward environments.
+SSE achieves faster convergence across tasks due to its early termination of episodes upon subgoal failure and a reduced number of transitions per episode.
+These features minimize unnecessary computation and accelerate learning.
+This advantage is particularly evident in long-horizon tasks such as AntKeyChest.
+In these tasks, conventional HRL methods consume many timesteps even during failed attempts and require extended horizons for high-level planning.
+In shorter-horizon scenarios like AntMazeBottleneck, the computational benefits of SSE are less significant since subgoal transitions and failure terminations occur less frequently.
+The operations introduced by SSE, including the evaluation of subgoal completion and uniform sampling, are lightweight and have linear time complexity.
+Consequently, SSE imposes minimal computational overhead while maintaining stable training dynamics.
 Table 4: Average per-episode computation time (in seconds) and the number of episodes required to reach 60% success rate.
@@
 To gain deeper insight into the behavior of SSE, we present trajectory analyses across representative environments.
-Figure 11: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeBottleneck across: (a) early stage, (b) mid training, and (c) task success.
 Fig.
 11 illustrates learning progression in AntMazeBottleneck.
-At the early stage (a), the agent has not yet discovered feasible subgoals, and the high-level policy $\pi^{h}$ selects subgoals largely at random.
+In the early stage (a), the agent has not yet discovered feasible subgoals, causing the high-level policy $\pi^{h}$ to select subgoals largely at random.
 During this phase, the exploration policy promotes coverage by gradually expanding into novel regions.
-In the mid stage (b), the agent fails to traverse the narrow bottleneck, primarily because the low-level policy $\pi^{l}$ lacks sufficient experience in that region and the associated failure statistics $\mathrm{ratio}_{\mathrm{fail}}$ remain underrepresented.
-By the final stage (c), SSE successfully leverages failure-aware refinement and subgoal selection, allowing $\pi^{h}$ to navigate through the bottleneck and reach the goal reliably.
-Figure 12: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeDoubleBottleneck across: (a) early stage, (b) mid training, and (c) task success.
-A similar trend is observed in AntMazeDoubleBottleneck as Fig.
+In the mid stage (b), the agent fails to traverse the narrow bottleneck, primarily because the low-level policy $\pi^{l}$ lacks sufficient experience in that region and the associated failure statistics $\rho_{\mathrm{f}}$ remain underrepresented.
+By the final stage (c), SSE successfully leverages failure-aware refinement and subgoal selection.
+This allows $\pi^{h}$ to navigate through the bottleneck and reach the goal reliably.
+Figure 11: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeBottleneck across: (a) early stage, (b) mid training, and (c) task success.
+A similar trend is observed in AntMazeDoubleBottleneck, as shown in Fig.
 12.
@@
 In the intermediate phase (b), the agent again struggles with the second bottleneck due to insufficient failure feedback.
-As training progresses (c), subgoals selected by $\pi^{h}$ become more effective, and the updated path refinement guides the agent through safer routes, enabling successful traversal of both bottlenecks and completion of the long-horizon task.
-Figure 13: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntKeyChest across: (a) early stage, (b) after reaching the key, and (c) reaching the goal with the key (task success).
+As training progresses (c), subgoals selected by $\pi^{h}$ become more effective.
+The updated path refinement guides the agent through safer routes, enabling successful traversal of both bottlenecks and completion of the long-horizon task.
+Figure 12: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntMazeDoubleBottleneck across: (a) early stage, (b) mid training, and (c) task success.
 In Fig.
@@
 In the final stage (c), the agent demonstrates the ability to sequentially reason over subtasks, first reaching the key and then navigating to the final goal with the required flag active, completing the task successfully.
+Figure 13: Trajectory analysis of SSE subgoals $\tilde{g}_{\max}$ in AntKeyChest across: (a) early stage, (b) after reaching the key, and (c) reaching the goal with the key (task success).
 E.2 Additional Ablation Studies
 Component Evaluation
-Figure 14: Component evaluation results in various maps
 Fig.
 14 presents the performance of SSE variants with key components ablated.
-Specifically, we evaluate the following variants: SSE without path refinement (PR), without $\tilde{g}_{\mathrm{novel}}$, $\tilde{g}_{\max}$, or goal $g$ in $\pi^{\mathrm{exp}}$, without the entire $\pi^{\mathrm{exp}}$, and with a fixed low-level step size (FLS).
-In complex tasks like AntKeyChest and AntDoubleKeyChest, removing max-value guidance ($\tilde{g}_{\max}$) or novelty-based exploration ($\tilde{g}_{\mathrm{novel}}$) leads to notable performance drops, due to poor goal-space coverage and delayed reward discovery.
-In contrast, simpler tasks like AntMazeBottleneck are less sensitive to these components.
-Path refinement similarly plays a larger role in long-horizon settings by mitigating unreliable transitions, while offering little benefit in short tasks with fewer failures.
-Removing the explicit goal signal (SSE w/o $g$) or exploration policy (SSE w/o $\pi^{\mathrm{exp}}$) consistently harms performance across all environments, confirming the importance of goal-conditioned reasoning and structured exploration.
-Lastly, replacing SSE’s flexible execution with a fixed-length scheduler (SSE w FLS) shows little effect in short tasks, but severely limits learning in long-horizon, sparse-reward environments like AntDoubleKeyChest.
+We evaluate four variants: (1) SSE w/o $\mathcal{B}_{F}^{h}$ which replaces FER with a standard replay buffer, (2) SSE with HER which uses a conventional fixed-step policy, (3) SSE w/o $\pi^{\mathrm{exp}}$ which disables decoupled exploration, and (4) SSE w/o PR which omits path refinement.
+Removing path refinement degrades performance in long-horizon settings where mitigating unreliable transitions is crucial.
+Similarly, disabling the exploration policy consistently impairs goal-space coverage across all environments.
+Critically, SSE with HER and SSE w/o $\mathcal{B}_{F}^{h}$ fail in complex tasks like AntDoubleKeyChest despite achieving minor success in simpler maps.
+This failure demonstrates that conventional HER offers insufficient feasibility signals.
+Furthermore, it confirms that FER is vital for providing stop-on-failure feedback to teach the high-level policy to avoid unreliable subgoals.
+These components collectively ensure the decisive signals required for long-horizon planning.
+Figure 14: Component evaluation results in various maps
 Effect of Path Refinement Scaling Factor $c_{\mathrm{dist}}$
+Fig.
+15 shows how the scaling factor $c_{\mathrm{dist}}$ affects performance.
+This parameter controls the trade-off between path efficiency and safety.
+In simpler tasks like AntMazeBottleneck, its impact is minimal.
+In more complex, long-horizon tasks, its role becomes more pronounced.
+A moderate value, such as $c_{\mathrm{dist}}=5$, provides an effective balance, guiding the planner away from failure-prone regions without being overly conservative.
+In contrast, excessively high values (e.g., 10 or 20) can lead to inefficient detours, while a value of 1 may not sufficiently penalize risky paths.
+Overall, the performance remains high for values between 2 and 10, indicating a wide effective range.
 Figure 15: Distance scail $c_{\mathrm{dist}}$ analysis in various maps
+Effect of Grid Size $d_{\mathcal{G}}$
 Fig.
-15 illustrates the effect of the path refinement scaling factor $c_{\mathrm{dist}}$ on performance across different environments.
-In short-horizon tasks such as AntMazeBottleneck, the sensitivity to $c_{\mathrm{dist}}$ is relatively low, and changes in its value have little impact on overall performance.
-In contrast, for longer-horizon tasks such as AntKeyChest and AntDoubleKeyChest, the effect of $c_{\mathrm{dist}}$ becomes more significant.
-When the value increases from 1 to 5, the agent benefits from more reliable path selection due to moderate penalty scaling, which helps avoid failure-prone regions while maintaining reachability.
-However, setting $c_{\mathrm{dist}}$ to overly large values such as 10 or 20 leads to frequent detours and unstable planning, ultimately degrading both learning efficiency and task success.
-Effect of Grid Size $d_{\mathcal{G}}$
+16 illustrates the effect of grid resolution $d_{\mathcal{G}}$.
+This parameter balances the granularity of novelty detection with the stability of failure statistics.
+The analysis shows that performance is not highly sensitive to this parameter within the tested range.
+For instance, in AntMaze environments, there is little significant difference in performance for $d_{\mathcal{G}}$ values of 1, 2, and 4.
+A coarse grid ($d_{\mathcal{G}}=4$) may merge distinct regions, while an overly fine grid ($d_{\mathcal{G}}=1$) can make failure statistics less reliable.
+Thus, we find that a moderate resolution ($d_{\mathcal{G}}=2$ for AntMaze) provides a suitable balance.
+The optimal choice is dependent on the environment’s scale, with smaller, structured spaces like Reacher benefiting from a finer grid ($d_{\mathcal{G}}=1$).
 Figure 16: Grid size $d_{\mathcal{G}}$ analysis in various maps
+Effect of Exploration Ratio $\eta$
 Fig.
-16 illustrates how the grid resolution $d_{\mathcal{G}}$, used to partition the goal space for novelty identification, affects exploration.
-A coarse grid with $d_{\mathcal{G}}=4$ fails to capture fine-grained novelty, limiting the ability to distinguish under-visited regions.
-This shortcoming is pronounced in environments with narrow passages, such as AntMazeBottleneck, where structural bottlenecks are poorly reflected.
-In contrast, $d_{\mathcal{G}}=1$ creates overly fine partitions, fragmenting the goal space and introducing noise into failure statistics, which leads to unstable path updates and slower learning.
-These results highlight the need for an appropriate balance between granularity and stability.
-Empirically, we use $d_{\mathcal{G}}=2$ for AntMaze environments to ensure reliable exploration while avoiding excessive updates.
-For Reacher environments, which feature a smaller and more structured goal space, a finer resolution of $d_{\mathcal{G}}=1$ is applied to retain sufficient detail without incurring instability.
-Effect of Exploration Ratio $\eta$
+17 shows how performance varies with the exploration ratio $\eta$, which controls the classic exploration-exploitation trade-off.
+In simple environments like AntMazeBottleneck, performance is consistent across a wide range of $\eta$ values.
+In long-horizon tasks like AntDoubleKeyChest that require discovering intermediate objectives, the influence of $\eta$ is naturally greater, but effective performance can be achieved with straightforward tuning.
+As shown in our experiments, setting the exploration ratio within a small range of 0.1 to 0.2 is sufficient to achieve strong performance across all tasks.
+This indicates that while the balance is important, extensive tuning of this hyperparameter is not required.
 Figure 17: Exploration ratio $\eta$ analysis in various maps
+Effect of Reachability Check Threshold $\lambda$
 Fig.
-17 presents the performance sensitivity with respect to the exploration ratio $\eta$.
-In short-horizon environments like AntMazeBottleneck, where high-level decisions are minimal and navigation is relatively simple, performance remains stable across a wide range of $\eta$ values.
-In such settings, large $\eta$ values are not necessary and may even hinder learning by reducing the frequency of value-guided subgoal selection, which slows the improvement of the high-level policy $\pi^{h}$.
-In contrast, long-horizon environments such as AntDoubleKeyChest, which require intermediate decision-making and structured exploration, exhibit strong sensitivity to $\eta$.
-A small $\eta$ impedes the discovery of intermediate goals, while a large $\eta$ restricts exploitation of the learned value function $Q^{h}$ after critical behaviors are found.
-These results underscore the need to carefully tune $\eta$ to maintain a proper balance between exploration and exploitation, particularly in tasks with extended temporal dependencies.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The abstract and introduction clearly state the proposed framework (SSE), its key features and highlight consistent improvements across benchmark tasks in Section 1.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Section 6 describes limitations in high-dimensional goal spaces and in representing tasks that require abstract skills not decomposable into spatial subgoals.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This work is empirical in nature and does not include formal theoretical claims that require proof.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: The proposed method is described in detail in Section 4, with full implementation details, including algorithmic procedures, training dynamics, and update rules, provided in Appendix B.2.
-These descriptions are sufficient to reproduce the main experimental results without requiring access to the source code.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24.
-Justification: We provide the full codebase and experiment scripts in the supplemental material, including detailed instructions for reproducing all main results.
-The code is organized for immediate execution with provided environment configurations.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: All relevant training and evaluation details, including optimizer configurations, hyperparameter settings, and scenario-specific parameter choices, are thoroughly documented in Appendix C.
-These descriptions are sufficient to fully interpret and reproduce the experimental results.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: All experimental results are averaged over 5 independent runs with different random seeds, and shaded regions in all figures represent the corresponding standard deviations, as detailed in Section 5.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: Details of the computational setup, including GPU types, memory, and per-run training time, are provided in Appendix C and Appendix D.2.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research is based on standard benchmark environments, poses no foreseeable societal or ethical risks, and adheres fully to the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Broader Impact is discussed in Appendix A.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper does not involve the release of models or data with potential for misuse or dual-use concerns.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All environments and baseline implementations are used under their official licenses and appropriately cited in the paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: All introduced methods and environments are provided with detailed documentation to support reproducibility and understanding.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The research does not involve any human subjects or crowdsourcing components.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The research does not involve human subjects or any procedures requiring IRB approval.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: No LLMs were used in the development of the core methodology; usage was limited to writing and editing support.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+18 illustrates the performance variation with respect to the reachability threshold $\lambda$.
+We employ the standard environment-provided value as the default.
+Performance remains stable as long as $\lambda$ is set smaller than the final task goal threshold, ensuring sufficient precision for intermediate waypoints.
+Conversely, extremely small values like 0.1 degrade performance in long-horizon tasks by forcing the low-level policy to achieve excessive precision.
+This hinders efficient exploration.
+These results confirm that the framework is robust to $\lambda$ and that the standard environment value offers an optimal balance.
+Figure 18: Distance threshold $\lambda$ analysis in various maps
+Effect of High-Level Gamma $\gamma^{h}$
+Fig.
+19 confirms that $\gamma^{h}=0.4$ yields optimal performance.
+Theoretically, SSE enforces strict, single-step reachability, which inherently condenses the high-level decision horizon.
+In this regime, a standard high discount factor (e.g., $\gamma^{h}=0.99$) fails to distinguish between immediate and delayed successes.
+Conversely, a lower $\gamma^{h}$ sharpens credit assignment, making the value function $Q^{h}$ highly sensitive to temporal inefficiencies.
+This sensitivity guides the policy to prioritize direct, high-probability transitions, effectively delineating the reachability frontier and accelerating convergence.
+Figure 19: High-level gamma $\gamma^{h}$ analysis in various map
+E.3 Comparison on Additional Environments
+To assess generalizability, we conducted experiments on AntPush, AntFall, and Ant4Rooms.
+Since conventional baselines like HIRO failed to learn in these settings, we focus our comparison on the graph-based methods BEAG and NGTE.
+As shown in Fig.
+20, standard graph-based algorithms struggle in tasks with complex contact dynamics or invisible obstacles.
+Their rigid reliance on shortest-path planning often leads to failure in environments like AntPush and AntFall.
+In contrast, SSE demonstrates superior performance.
+We attribute this to the decoupled exploration policy $\pi^{\mathrm{exp}}$.
+This mechanism allows the high-level policy $\pi^{h}$ to leverage diverse exploration data beyond the graph’s shortest path, facilitating robust planning under complex physical constraints.
+Note that in AntPush, we froze graph edge updates after initial convergence to mitigate instability from object interactions.
+SSE maintained its performance advantage even under this stabilization protocol.
+Figure 20: Performance comparison on additional environments
+E.4 More Random Seeds
+To validate the statistical stability of our results, we supplemented our initial experiments (run with 5 seeds, a common practice in this field) by re-running the key complex tasks with 10 random seeds.
+We focused on tasks like AntKeyChest and AntDoubleKeyChest (Fig.
+6, Fig.
+8), where the standard deviation range of SSE with 5 seeds already showed a clear separation from the baselines.
+As shown in Fig.
+21, the experiment with 10 seeds produces a nearly identical performance curve, and no significant change in the mean success rate was observed compared to 5 seeds.
+More importantly, we re-affirmed that the standard deviation bounds remained clearly separated from the baseline algorithms.
+This confirms that the superior performance of SSE was not achieved by chance with a low seed count and that our findings are reliable and statistically stable.
+Figure 21: Performance comparison of SSE with 10 seeds